### v4

# references.bib

Abstract
This research comprehensively describes the basic theory of transversally Heisenberg elliptic operators, and investigates the index theory of Heisenberg elliptic and transversally Heisenberg elliptic operators from the perspective of $KK$-theory, applying Kasparov’s methodology. Moreover, the analysis methodically examines specific conditions, with a focus on the Fourier transform of the nilpotent group $C^{\ast}$-algebra. We demonstrate enhanced methods for analyzing the hypoellipticity of operators, presenting a robust framework for defining and understanding transversal Heisenberg ellipticity in a $KK$-theoretic context. This work provides a solid foundation for future research into the properties of differential operators in complicated manifolds.
Keywords. Heisenberg elliptic, transversally Heisenberg elliptic, index theory, $KK$-theory
Dedication
This thesis is dedicated to my parents, whose unwavering support and encouragement have been my guiding lights throughout this academic journey. To my friends, your camaraderie and understanding made the challenges more manageable and the successes more joyful.
Declaration
I hereby declare that this thesis is entirely my own work. I affirm the following:
  • (a)
    This work has not been submitted for any other degree or academic qualification.
  • (b)
    All sources of information and ideas used in this thesis are properly acknowledged.
  • (c)
    Contributions made by others in the form of ideas, guidance, or any other support are duly recognized and appropriately cited.
  • (d)
    The research presented in this thesis adheres to the ethical standards and guidelines of academic integrity.
Acknowledgements
I would like to express my sincere gratitude to my supervisor, Professor Kato Tsuyoshi, for his invaluable ideas and guidance for this research. Professor Gennadi Kasparov’s patient leadership and Professor Nigel Higson’s valuable advice were instrumental in shaping this research.
Special thanks to the Graduate Education Support Organization Program / Science and Technology Innovation Creation Fellowship for their generous support, enhancing the depth and scope of this work.
Introduction
Let $M$ be a smooth and closed manifold, $E,F\to M$ be vector bundles, and $P:C^{\infty}(M;E)\to C^{\infty}(M;F)$ be a differential operator over $M$. For elliptic operators, the spaces $\text{Ker}(P)$ and $\text{Ker}(P^{\ast})$ are both finite-dimensional, and the analytical index is defined as $\text{Ind}_{a}(P)=\dim\text{Ker}(P)-\dim\text{Coker}(P)\in\mathbb{Z}$. When a Lie group $G$ acts on $M$ and the operator $P$ is elliptic, the spaces $\text{Ker}(P)$ and $\text{Coker}(P)$ can be viewed as elements of $R(G)$, with the analytical index defined as $\text{Ind}_{a}(P)=[\text{Ker}(P)]-[\text{Coker}(P)]\in R(G)$. The Atiyah-Singer index theorem [atiyah1968index2][atiyah1968index1] states that in each case, the analytical index defined above is equal to the topological index, which only depends on the topological characteristics of $P$ and $M$.
For the case of transversally elliptic operators, Atiyah [atiyah2006elliptic] introduced the idea of defining the analytical index as a distribution over the Lie group $G$. Berline and Vergne [berline1996chern] introduced a cohomological index formula, equating the analytical index to the cohomological index for transversally elliptic operators.
Kasparov [kasparov2017elliptic][kasparov2022k] generalized it for locally compact manifolds under properly supported Lie group actions, framing a comprehensive theory on the index of elliptic and transversally elliptic operators through $KK$-theory. In this theory both the symbol class $[\sigma(P)]$ and the index class $[P]$ are represented by $KK$-classes, providing the index theorem through $KK$-products.
The study of index theory for hypoelliptic operators is challenging due to the simple condition. Additional conditions are required to define the analytical index accurately.
Boutet de Monvel [de1978index] gave the topological formula for the analytical index of Toeplitz operators. van Erp [van2010atiyah1][van2010atiyah2] generalized it to arbitrary contact manifolds, based on the work by Epstein and Melrose [epstein2004heisenberg], for Heisenberg elliptic operators without vector bundle coefficients. Later van Erp and Baum [baum2014k] generalized van Erp’s formula to allow vector bundle coefficients on contact manifolds. Their formula allows us to compute the index of some differential operators of Hörmander’s sum of squares of rank $2$. Androulidakis, Mohsen, and Yuncken [androulidakis2022pseudodifferential] generalized the classical regularity theorem of elliptic operators to maximally hypoelliptic differential operators, and Mohsen[mohsen2022index] established an index formula for $\ast$-maximally hypoelliptic differential operators on closed manifolds.
On a closed manifold with filtration, or equivalently on a closed manifold endowed with a regular foliation, the index theorem for Heisenberg elliptic (later on, H-elliptic) operators can be reduced as a special case of Mohsen’s maximally hypoelliptic index theorem. In the regular foliation case, the characteristic set defined in [androulidakis2022pseudodifferential] is the entire groupoid $Gr(H)$. Hence, the $\ast$-maximally hypoelliptic condition is consistent with the H-ellipticity, which requires for each nontrivial unitary irreducible representation $\pi$, the action on the H-cosymbol $\sigma_{H}(P,x,\pi)$ defined by the filtration is bijective. We will also provide an index theorem for H-elliptic operators from the $KK$-theoretic viewpoint in Section 3.2.
This research introduces transversally H-elliptic operators on filtered manifolds under a $G$-action, a concept that is inspired by Kasparov’s tangent-Clifford symbol class $[\sigma^{tcl}(P)]$ in transversally elliptic index theory [kasparov2022k]. Transversal H-ellipticity requires the Rockland condition’s fulfillment in conjunction with a leafwise elliptic symbol class. Concretely, We define that:
Definition 0.0.1.
Let $M$ be a compact filtered Riemannian manifold and $G$ be a Lie group acting on $M$ isometrically and preserving the filtered structure. A pseudo-differential operator $P$ of 00-order is called transversally H-elliptic if at each point $x\in M$, the multiplier term
$(1-\sigma_{H}^{2}(P,x))(1+\sigma_{H}(\Delta_{G},x))^{-1}\in M(E_{x}\otimes_{\mathbb{C}}C^{\ast}(Gr(H)_{x})),$
as an operator on $C^{\ast}(Gr(H)_{x})$-module is compact, where $\Delta_{G}$ is the sub-Laplacian operator in $G$-orbit directions.
In Chapter 4, We will prove that Definition 0.0.1 produces an appropriate tangent-Clifford symbol class $[\sigma^{tcl}_{H}(P)]$ and an index class $[P]$ for each transversally H-elliptic operator $P$. When $P_{0}$ is a transversally H-elliptic differential operator, let $P$ be the bounded transform of $P_{0}$. Then, there is an index formula in $KK$-homology in which $[\sigma^{tcl}_{H}(P)]$ and $[P]$ can be related using $KK$-products. Here is the main theorem of the paper:
Theorem 0.0.2.
Let $M$ be a closed filtered Riemannian manifold and $G$ be a Lie group isometrically acting on $M$ and preserving the filtered structure. Let $P_{0}$ be a $G$-invariant differential operator on $M$ with its bounded transform
$P=\begin{pmatrix}0&P_{0}^{\ast}/(1+P_{0}P_{0}^{\ast})\\ P_{0}/(1+P_{0}^{\ast}P_{0})&0\end{pmatrix}$
being transversally H-elliptic of order 00. Then, we have the following formula
$[P]=j^{G}([\sigma^{tcl}_{H}(P)])\otimes_{C^{\ast}(G,C^{\ast}(Gr(H)))}j^{G}([\mathcal{E}])\otimes_{C^{\ast}(G,C_{0}(T^{\ast}M))}[\mathcal{D}_{M,\tilde{\Gamma}}^{cl}],$
where $[\sigma^{tcl}_{H}(P)]\in K_{0}^{G}(C^{\ast}(Gr(H))\otimes_{C(M)}Cl_{\tilde{\Gamma}}(M))$ is the tangent-Clifford H-cosymbol for $P$, $[\mathcal{E}]\in KK_{G}(C^{\ast}(Gr(H)),C_{0}(T^{\ast}M))$ refers to the Connes-Thom class and $[\mathcal{D}_{M,\tilde{\Gamma}}^{cl}]\in K^{0}(C^{\ast}(G,$ $Cl_{\tilde{\Gamma}}(TM)))$ indicates the Clifford–Dolbeault element.
Determining whether an arbitrary pseudo-differential operator meets the criteria outlined in our defined theory hinges on the Fourier transform theory for nilpotent group $C^{\ast}$-algebras. For an element $\phi\in C^{\ast}(\mathcal{G})$, where $\mathcal{G}$ represents a nilpotent Lie group, its representations $\pi(\phi)$ constitute an operator-valued function on the spectrum $\widehat{\mathcal{G}}$, with some specific conditions. This theory extends to the Fourier transform of multiplier elements of the group $C^{\ast}$-algebra, transforming a multiplier of $C^{\ast}(\mathcal{G})$ into an operator family parametrized by $\widehat{\mathcal{G}}$. As a result, a multiplier of $C^{\ast}(\mathcal{G})$ is Fourier-transformed to a family of operators parametrized by the representation space $\widehat{\mathcal{G}}$, and the properties of this operator family reflect if the multiplier is an element of $C^{\ast}(\mathcal{G})$ itself or not. Explicitly, for an element $e\in M(C^{\ast}(\mathcal{G}))$, if we can verify that the Fourier transform $\mathcal{F}(e)\in\mathcal{F}(M(C^{\ast}(\mathcal{G})))$ actually lies in $\mathcal{F}(C^{\ast}(\mathcal{G}))$, then $e\in\mathcal{K}(C^{\ast}(\mathcal{G}))=C^{\ast}(\mathcal{G})$. See the following diagram:
[…]
Take $\mathcal{G}=Gr(H)$, we will employ this method to verify the compactness of the $C^{\ast}(Gr(H)_{x})$-module multipliers:
Theorem 0.0.3.
For each $x\in M$, suppose that $Q$ is a multiplier of the $C^{\ast}(Gr(H)_{x})$ module, $Q\in M(E_{x}\otimes C^{\ast}(Gr(H)_{x}))$. Then, $Q$ is compact , i.e.,
$Q\in\text{End}(E_{x})\otimes C^{\ast}(Gr(H))_{x}$
if and only if the operator field defined by
$\phi=\{\pi_{\gamma}(Q)\}_{\gamma\in\widehat{Gr(H)}_{x}}\subseteq\text{End}(E_{x})\otimes\mathit{l}^{\infty}(\widehat{Gr(H)}_{x})$
lies in the sub-algebra $\text{End}(E_{x})\otimes B^{\ast}(\widehat{Gr(H)}_{x})$. The algebra $B^{\ast}(S)$ is defined in Definition 6.1.5.
An illustrative example of transversally H-elliptic operators is given by the operator $P_{\gamma}=\Delta_{n-k}+i\gamma T$ on the Heisenberg group $M=\mathbb{H}_{n}$ with transition action by $\mathbb{R}^{2k}$ on the first $2k$-coordinates. Here, $\Delta_{n-k}=-\sum_{i=k+1}^{n}({X_{i}}^{2}+Y_{i}^{2})$ means the sub-Laplacian operator on the sub-manifold $M^{\prime}\cong\mathbb{H}_{n-k}$ identified by coordinates $x_{k+1},y_{k+1},\dots,x_{n},y_{n},z$. The function $\gamma$ on $M$ is assigned such that $P_{\gamma}$ on $M^{\prime}$ is H-elliptic, see [baum2014k]. Utilizing Theorem 0.0.3, we can show that $P_{\gamma}$ is transversal H-elliptic on $M$.
Structure of the paper
  • Chapter 1 sets the stage with preliminaries on symbolic calculus on filtered manifolds and Kirillov’s orbit method for Heisenberg calculus of differential and pseudo-differential operators. It also touches upon Connes’ [connes1994noncommutative] groupoid approach for H-elliptic cases.
  • Chapter 2 introduces the $KK$-theoretic framework of the index theorem for elliptic and transversally elliptic operators, as detailed in [kasparov2022k][kasparov2017elliptic].
  • Chapter 3 articulates the general index theorem for H-elliptic operators, extending van Erp’s findings on contact manifolds [van2010atiyah1] and presenting the theorem from a $KK$-theoretic perspective.
  • Chapter 4 gives a definition for transversally H-elliptic operators, establishing the framework for defining symbol and index classes.
  • Chapter 5 is dedicated to the main index theorem for transversally H-elliptic operators. We give a proof for the H-elliptic index theorem in Chapter 3 and extend it to transversally H-elliptic operators.
  • Chapter 6 introduces a Fourier transform theory for $C^{\ast}$-algebra of nilpotent Lie groupoids [beltictua2017fourier] to verify the compactness condition of Definition 0.0.1.
  • In the Appendix, we exhibit other possible conditions to define the transversally H-elliptic operators. We include prototypes of these definitions, from the definitions or certain crucial properties of transversally elliptic or H-elliptic operators.
Notations
  • We denote by $G$ the Lie group act on manifold $M$, by $\mathcal{G}$ a groupoid or a bundle of nilpotent Lie groups over the manifold $M$.
  • We write $H^{\bullet}$ to mean the filtration structure over the manifold $M$, and $\mathcal{H}$ or $\mathcal{H}_{\pi}$ for the representation space of some unitary irreducible representation of the Lie group $\mathcal{G}\to U(\mathcal{H}_{\pi})$.
  • We write $\sigma_{H}(P)$ or $\sigma_{H}(P,x)$ to mean the H-cosymbol of $P$ and $\sigma(P,x,\pi)$ to mean the representation $\pi$ on H-cosymbol $\pi(\sigma_{H}(P,x))$.
  • If $E$ is a vector bundle over $M$, we will denote by $C^{\infty}(M;E)$ the space of smooth sections of $E$, $C(M;E)$ the space of continuous sections of $E$ and $C_{0}(M;E)$ the space of continuous sections of $E$ vanishing at infinity. We denote $C_{0}(E)$ to mean the space of continuous functions on $E$ (a topological space) vanishing at infinity.
  • On the tangent bundle $TM$, we denote its coordinate by $(x,\xi)$. Its tangent space $T(TM)$ has coordinate $(x,\xi,\chi,\zeta)$, where $\chi$ is the differential coordinate with respect to $x$, and $\zeta$ is the differential coordinate with respect to $\xi$.
  • We denote by $\otimes_{D}$ the Kasparov product
    $\otimes_{D}:KK(A_{1},B_{1}\hat{\otimes}D)\times KK(D\hat{\otimes}A_{2},B_{2})\to KK(A_{1}\hat{\otimes}A_{2},B_{1}\hat{\otimes}B_{2}).$
  • We denote by $M(A)$ the multiplier algebra for a $C^{\ast}$-algebra $A$.
Chapter 1 Preliminaries for H-elliptic theory
In this chapter, we introduce fundamental concepts related to filtered manifolds and the symbolic calculus of operators on these manifolds.
1.1 Filtered manifold and the osculating groupoid
Let $M$ be a smooth closed manifold with a filtration of the tangent bundle
$0=H^{0}\subseteq H^{1}\subseteq H^{2}\subseteq\cdots\subseteq H^{n-1}\subseteq H^{n}=TM$
by a sequence of vector bundles $H^{i}$, satisfying
$[C^{\infty}(M;H^{i}),C^{\infty}(M;H^{j})]\subseteq C^{\infty}(M;H^{i+j}),\ \forall 1\leq i,j\leq n.$
Here, we follow the convention that $H^{i}=TM$ for $i>n$.
Assume that $M$ is equipped with a proper isometric action by a locally compact group $G$.
Definition 1.1.1.
The filtration structure $H^{\bullet}$ is $G$-invariant if for each $g\in G,x\in M$, the differential of the group action $g:M\to M,x\mapsto g(x)$,
$dg:T_{x}M\to T_{g(x)}M$
maps $H^{i}_{x}$ to $H^{i}_{g(x)}$ for all $0\leq i\leq n$.
We will always consider $G$-invariant filtrations in the following contents.
Lemma 1.1.2.
At each point $x$ on the filtered manifold $M$, there is a simply-connected nilpotent Lie group structure on $T_{x}M$, called the osculating group at $x$.
Proof.
Define the graded bundle over $M$,
$gr(H):=\bigoplus_{i}H^{i}/H^{i-1}.$
At each point $x\in M$, $gr(H)_{x}=H^{i}_{x}/H^{i-1}_{x}$ has a graded nilpotent Lie algebra structure induced from
$[X(x),Y(x)]:=[X,Y](x)\ \text{mod}\ H^{i+j-1}_{x},X\in H^{i},Y\in H^{j}.$
It can be proved that $[X,Y](x)$ mod $H^{i+j-1}_{x}$ depends only on $X(x)$ mod $H^{i-1}_{x}$ and $Y(x)$ mod $H^{j-1}_{x}$. Therefore, the Lie bracket structure over $gr(H)_{x}$ is well defined for each $x\in M$. The graded algebra $gr(H)_{x}$ at each point is called the osculating Lie algebra. Then there is a corresponding simply-connected nilpotent Lie group for each $x$, we denote it by $Gr(H)_{x}$. ∎
Definition 1.1.3.
The Lie groupoid
$Gr(H):=\bigsqcup_{x\in M}Gr(H)_{x}$
is called the osculating Lie groupoid of $M$ with filtration structure $H^{\bullet}$.
The osculating groupoid serves as the non-commutative counterpart of the tangent bundle $TM$.
Remark 1.1.4.
The graded Lie algebra possesses a natural group of dilation automorphisms. For any $\lambda>0$, let $\delta_{\lambda}\in\text{Aut}(gr(H))$ denote the bundle automorphism given by multiplication with $\lambda^{i}$ on the grading component $gr(H)^{i}$. For each $x\in M$, this restricts to an automorphism $\delta_{\lambda,x}\in\text{Aut}(gr(H)_{x})$ of the osculating Lie algebra such that $\lim_{\lambda\to 0}\delta_{\lambda,x}=0$. Evidently, $\delta_{\lambda_{1}\lambda_{2}}=\delta_{\lambda_{1}}\delta_{\lambda_{2}}$ for all $\lambda_{1},\lambda_{2}>0$.
1.2 Kirillov’s orbit method
The primary goal of Kirillov’s orbit method, as outlined in [kirillov2004lectures], is to characterize the unitary dual $\widehat{G}$ as a topological space and to construct the unitary irreducible representations $\pi_{\Omega}$ associated with a coadjoint orbit $\Omega\in\mathfrak{g}^{\ast}$.
Definition 1.2.1.
A coadjoint orbit is an orbit of a Lie group $G$ in the space $\mathfrak{g}^{\ast}$ dual to $\mathfrak{g}=\text{Lie}(G)$, on which the Lie group $G$ acts through the coadjoint representation.
The unitary irreducible representation $\pi_{\Omega}$ associated with each orbit $\Omega\subseteq\mathfrak{g}^{\ast}$ is obtained as follows:
  • Select any point $F\in\Omega$;
  • Identify a subalgebra $\mathfrak{h}$ of maximal dimension that is subordinate to $F$, i.e. $F|_{[\mathfrak{h},\mathfrak{h}]}=0$;
  • Define a subgroup $H=\exp\mathfrak{h}$ and construct the $1$-dimensional unitary irreducible $\rho_{F,H}$ of $H$ using the formula
    $\rho_{F,H}(\exp(X))=e^{2\pi i\langle F,X\rangle};$
  • $\pi_{\Omega}=\text{Ind}^{G}_{H}(\rho_{F,H})$.
Kirillov’s orbit method asserts the following.
Theorem 1.2.2 ([kirillov2004lectures]).
This map from coadjoint orbits $\Omega\in\mathcal{O}(G)$ in $\mathfrak{g}^{\ast}$ to unitary irreducible representations of $\pi_{\Omega}\in\widehat{G}$ establishes a one-to-one correspondence.
The topology of $\widehat{G}$ is defined by considering the space $\mathcal{O}(G)$ of coadjoint orbits as the quotient space of $\widehat{G}$ with the quotient topology.
Example 1.2.3.
In the case of Heisenberg group $\mathbb{H}_{n}\cong\mathbb{R}^{2n+1}$, the family of representations, corresponding to all co-adjoint orbits, can be described as
  • 1.
    of the form $\lambda T$ for $\lambda\in\mathbb{R}-\{0\}$, associated with the Schrödinger representation $\pi_{\lambda}$.
    $\pi_{\lambda}(x,y,t)\xi(s):=e^{-2\pi i\lambda-2\pi i\frac{\lambda}{2}x\cdot y+2\pi i\lambda s\cdot y}\xi(s-x),$
    for $s\in\mathbb{R}^{n}$, $\xi\in L^{2}(\mathbb{R}^{n})$, $(x,y,t)\in\mathbb{H}_{n}$. The infinitesimal operator of $\pi_{\lambda}$ can be written as follows:
    $d\pi_{\lambda}(X_{j})=i\lambda s_{j},d\pi_{\lambda}(Y_{j})=\frac{\partial}{\partial s_{j}},d\pi_{\lambda}(T)=i\lambda,$
    where
    $X_{j}=\frac{\partial}{\partial x_{j}}+\frac{1}{2}ty_{j}\frac{\partial}{\partial t},Y_{j}=\frac{\partial}{\partial y_{j}}-\frac{1}{2}tx_{j}\frac{\partial}{\partial t},T=\frac{\partial}{\partial t},j=1,\dots,n.$
  • 2.
    of the form $\sum_{j=1}^{n}(x_{j}^{\prime}X_{j}^{\prime}+y_{j}^{\prime}Y_{j}^{\prime})$ where $x_{j}^{\prime},y_{j}^{\prime}\in\mathbb{R}$, associated with the $1$-dimensional representation
    $(x,y,t)\mapsto\exp(i(x\cdot x^{\prime}+y\cdot y^{\prime})),$
    where $x\cdot x^{\prime}$ and $y\cdot y^{\prime}$ denote the canonical scalar product on $\mathbb{R}^{n}$.
1.3 Groupoid $C^{\ast}$-algebra
Definition 1.3.1.
  • A groupoid is a set $\mathcal{G}$ together with:
    • (a)
      a distinguished subset $\mathcal{G}^{(0)}\subset\mathcal{G}$;
    • (b)
      sourse and range maps $r,s:\mathcal{G}\rightrightarrows\mathcal{G}^{(0)}$;
    • (c)
      a multiplication map $\mathcal{G}^{(2)}\to\mathcal{G}$, $(a,b)\to ab$, where
      $\mathcal{G}^{(2)}:=\{(a,b)\in\mathcal{G}\times\mathcal{G}\ |\ s(a)=r(b)\};$
    • (d)
      an inverse map $\mathcal{G}\to\mathcal{G}$, $a\to a^{-1}$, such that:
      • 1.
        $r(x)=x=s(x)$ for all $x\in\mathcal{G}^{(0)}$.
      • 2.
        $r(a)a=a=as(a)$ for all $a\in\mathcal{G}$.
      • 3.
        $r(a^{-1})=s(a)$ for all $a\in\mathcal{G}$.
      • 4.
        $s(a)=a^{-1}a$ and $r(a)=aa^{-1}$ for all $a\in\mathcal{G}$.
      • 5.
        $r(ab)=r(a)$ and $s(ab)=s(b)$ for all $(a,b)\in\mathcal{G}^{(2)}$.
      • 6.
        $(ab)c=a(bc)$ whenever $s(a)=r(b)$ and $s(b)=r(c)$.
  • A Lie groupoid is a groupoid $\mathcal{G}$ that is also a smooth manifold, such that the space of units $\mathcal{G}^{(0)}$ is a smooth submanifold of $\mathcal{G}$; source and range maps $s,r:\mathcal{G}\rightarrow\mathcal{G}^{(0)}$ are smooth submersions; multiplication is a smooth map $\mathcal{G}^{(2)}\rightarrow\mathcal{G}$; and the inverse map $x\mapsto x^{-1}$ is smooth $\mathcal{G}\rightarrow\mathcal{G}$. We write $\mathcal{G}^{x}=r^{-1}({x})$ and $\mathcal{G}_{x}=s^{-1}({x})$ for the range and source fibers at $x\in\mathcal{G}^{(0)}$, which are submanifolds of $\mathcal{G}$.
Definition 1.3.2.
  • A Haar system on a Lie groupoid $\mathcal{G}$ is a family of smooth measures $\{\mu_{x}\}$ supported on fibers $\mathcal{G}^{x}$ of $\mathcal{G}$ such that:
    • (a)
      For any $f\in C_{c}(\mathcal{G})$, the function
[…]
      is continuous and compactly supported;
    • (b)
      For all $\eta\in\mathcal{G}$,
[…]
  • Let $f,g\in C_{c}(\mathcal{G})$, define the convolution and involution maps:
[…]
    $f^{\ast}(\gamma):=\overline{f(\gamma^{-1})}.$
Lemma 1.3.3 ([renault2006groupoid]).
The operations above turn $C_{c}(\mathcal{G})$ into a $\ast$-algebra.
Definition 1.3.4.
Let $\mu=\{\mu_{x}\in G^{0)}\}$ be a Haar system on the Lie groupoid $\mathcal{G}$. The groupoid $C^{\ast}$-algebra $C^{\ast}(\mathcal{G})$ is defined to be the completion of the algebra $C_{c}(\mathcal{G})$ for the norm
$\lVert f\rVert:=\sup_{\pi}\ \lVert\pi(f)\rVert,$
where $\pi$ runs over all non-degenerate representations of $C_{c}(\mathcal{G})$.
1.4 Heisenberg calculus for differential operators and H-elliptic operators
Let $E,F\rightarrow M$ be the vector bundles, $P:C^{\infty}(M;E)\rightarrow C^{\infty}(M;F)$ is a differential operator on $M$.
Definition 1.4.1.
$P$ is of H-order $\leq k$ if $P$ can be locally expressed as a sum of monomials of the form
$\nabla_{X_{1}}\cdots\nabla_{X_{s}}\Phi,$
where $\Phi\in C^{\infty}(M;End(E,F))$, $\nabla$ represents a connection on $F$, $X_{i}\in C^{\infty}(M;H^{a_{i}})$ and $\sum_{i=1}^{s}a_{i}\leq k$. We denote by $\text{Diff}^{k}_{H}(M,E,F)$ the space of differential operators of H-order $\leq k$.
Definition 1.4.2.
Let $\pi:Gr(H)_{x}\rightarrow U(\mathcal{H}_{\pi})$ be an irreducible unitary representation. If $P\in\text{Diff}^{k}_{H}(M,E,F)$ can be locally expressed as the sum of monomials of the form $\nabla_{X_{1}}\cdots\nabla_{X_{s}}\Phi$ near $x$ with the maximal order equal to $k$, then the Heisenberg principal cosymbol of $P$ at $\pi$ can be defined as
$\sigma_{H}^{k}(P,x,\pi)=\sum_{\text{order}=k}d\pi([X_{1}]_{x})\cdots d\pi([X_{s}]_{x})\otimes\Phi_{x}:\mathcal{H}^{\infty}_{\pi}\otimes E_{x}\to\mathcal{H}^{\infty}_{\pi}\otimes F_{x},$
where
  • $d\pi:gr(H)_{x}\to B(\mathcal{H}^{\infty}_{\pi})$, $d\pi(X)=\frac{d}{dt}|_{t=0}\pi(\exp(tX))v$ is the differential of the representation $\pi$,
  • $[X_{i}]_{x}^{a_{i}}\in gr(H)_{x}$ is the class of $X_{i}\in H^{a_{i}}$ in $gr(H)_{x}$.
[…]
Remark 1.4.3.
  • According to Kirillov’s orbit method, each point $\xi\in gr(H)_{x}$ corresponds to a representation $\pi_{\xi}$ of $Gr(H)_{x}$, indicating that $\sigma(P,x,\pi_{\xi})$ parallels the evaluation of the ordinary symbol $\sigma(P)$ on points $(x,\xi)\in T^{\ast}M$.
  • The $KK$-classes we will define later and the resulting map $\sigma_{H}^{k}(P,x,\pi)$ depend only on the principal component of $P$, so we can select the bundle map $TM\to gr(H)$ to be the simplest one which maps $v\in H^{i}_{x}$ to $[v]\in H^{i}_{x}/H^{i-1}_{x}\subset gr(H)_{x}$.
Lemma 1.4.4 ([fischer2016quantization]).
The differential representation $d\pi$ satisfies
$\forall Y_{1},Y_{2}\in gr(H)_{x},\ d\pi(Y_{1})d\pi(Y_{2})-d\pi(Y_{2})d\pi(Y_{1})-d\pi([Y_{1},Y_{2}])=0.$
Thus, $d\pi$ can be extended to be defined on the universal enveloping algebra of $gr(H)_{x}$ for each $x\in M$.
Definition 1.4.5.
Let $P=\nabla_{X_{1}}\cdots\nabla_{X_{s}}\Phi$, then
  • At each point $x\in M$, we define
    $P_{x}=\sum_{\text{order}=k}([X_{1}]_{x})\cdots([X_{s}]_{x})\otimes\Phi_{x}\in\mathcal{U}(gr(H)_{x})\otimes\text{End}(E_{x},F_{x})$
    as the freezing coefficient of $P$ at $x$,
  • The H-cosymbol of $P$ is defined to be the family of operators
    $\sigma_{H}(P):=\{\sigma_{H}(P,x)\}_{x\in M}=\{P_{x}\}_{x\in M}.$
    on $\{Gr(H)_{x}\}$ parameterized by $x\in M$.
The Heisenberg principal cosymbol of $P$ at $\pi$ $\sigma_{H}^{k}(P,x,\pi)$ can be calculated by
$\sigma_{H}^{k}(P,x,\pi)=(d\pi\otimes 1)(P_{x}):\mathcal{H}^{\infty}_{\pi}\otimes E_{x}\to\mathcal{H}^{\infty}_{\pi}\otimes F_{x}.$
Remark 1.4.6.
$\sigma_{H}(P)$ defined above is different from the symbol in the ordinary sense. For an ordinary differential operator of order $m$, $P=\sum_{|\alpha|\leq m}a_{\alpha}\partial_{x}^{\alpha}$, freezing the coefficient at a point $x\in M$ induces the constant coefficient operator $P_{x}=\sum_{|\alpha|=m}a_{\alpha}(x)\partial_{\xi}^{\alpha}$ on $T_{x}M$. This family $\{P_{x}\}_{x\in M}$ is called the ordinary principal “cosymbol” of $P$, with its Fourier transform being the principal symbol of $P$ in the ordinary sense:
$\sigma(P)=\sum_{\lvert\alpha\rvert=m}a_{\alpha}(x)(i\xi)^{\alpha}.$
The principal cosymbol is Fourier transformed to the principal symbol of $P$, in which the Fourier transform can be regarded as the representation by elements in $T_{x}^{\ast}M$ at each point. However, in the filtered manifold cases, $C^{\ast}(Gr(H))$ is not commutative so the Fourier transform is hard to calculate directly. That is why we only defined the Heisenberg cosymbol but not the Heisenberg symbol. The Fourier transform of the H-cosymbols will be discussed in Chapter 6.
The H-cosymbol will be frequently used within the Heisenberg calculus, whereas the terms “cosymbol” and “symbol” will refer to the family of operators of freezing coefficients and its Fourier transform, respectively, in a conventional context.
Let $N=\dim Gr(H)$. Let $\{X_{j}\}^{N}_{j=1}$ represent a local basis of the vector bundle $gr(H)$ near $x\in M$. then, $\{Y_{j}=X_{j}(x)\}^{N}_{j=1}$ will be a basis of $gr(H)_{x}$ at each point near $x$. Each $Y_{j}\in gr(H)_{x}$ can be regarded as a first-order left-invariant differential operator on $Gr(H)_{x}$, the formal sum
$\sum Y_{1}^{\alpha_{1}}\dots Y_{N}^{\alpha_{N}}$
gives rise to a higher order left-invariant differential operator on $Gr(H)_{x}$. In fact, we have the following theorem:
Theorem 1.4.7 (Poincaré-Birkhoff-Witt theorem, [bourbaki1989lie], Ch 1, Sec.2.7).
Considering $\{X_{j}\}^{n}_{j=1}$ as a basis of the Lie algebra $\mathfrak{g}$ of nilpotent Lie group $\mathcal{G}$, each vector $X\in\mathfrak{g}$ can be regarded as a left-invariant differential operator on $C^{\infty}(G)$ defined as
$Xf(x):=\frac{d}{dt}|_{t=0}f(x\cdot\exp_{G}(tX)).$
Conversely, any left-invariant differential operator $T$ on $G$ can be uniquely expressed as a finite sum
$T=\sum c_{\alpha}X^{\alpha}$
for some multi-index $\alpha=(\alpha_{1},\dots,\alpha_{n})$, where all but a finite number of the coefficients $c_{\alpha}\in\mathbb{C}$ are zero. This establishes an identification between the universal enveloping algebra $\mathcal{U}(\mathfrak{g})$ and the space of the left-invariant differential operators on $\mathcal{G}$.
The following analysis illustrates how a representation $\pi$ interacts directly with the complete cosymbol class $\sigma_{H}(P)=\{\sigma_{H}(P,x)\}_{x\in M}$, without using the differential representation $d\pi$.
Let $\pi:Gr(H)_{x}\to U(\mathcal{H_{\pi}})$ be a unitary representation, we can construct
$\pi(\sigma_{H}(P,x)):E_{x}\otimes\mathcal{H}_{\pi}\to F_{x}\otimes\mathcal{H}_{\pi}.$
as follows:
By Theorem 1.4.7, $\mathcal{U}(gr(H)_{x})$ elements can be identified with left-invariant differential operators on $Gr(H)_{x}$. So the cosymbol $\sigma_{H}(P)$, which is a family of elements in $\mathcal{U}(gr(H)_{x})$ parametrized by $x$, can be considered a collection of left-invariant differential operators $\{\sigma_{H}(P,x)\}_{x\in M}$ on the family of manifolds $\{Gr(H)_{x}\}_{x\in M}$.
For each $x\in M$, since $\sigma_{H}(P,x)$ corresponds to a left-invariant differential operator on $Gr(H)_{x}$, it follows that for $f,g\in C_{c}^{\infty}(Gr(H)_{x})$, $P(f\ast g)=Pf\ast g$. This principle allows us to extend the operator to the group $C^{\ast}$-algebra $C^{\ast}(Gr(H)_{x})$ and yields
$\overline{\sigma_{H}(P,x)}:\text{Dom}(\overline{P})\subseteq E_{x}\otimes C^{\ast}(Gr(H)_{x})\to F_{x}\otimes C^{\ast}(Gr(H)_{x}).$
This can be considered to be a morphism between $C^{\ast}(Gr(H)_{x})$-Hilbert modules. Thus, the cosymbol $\sigma_{H}(P)=\{\sigma_{H}(P,x)\}$ parametrized by $x\in M$ can be extended to an unbounded multiplier of $C^{\ast}({Gr(H)})$
$\sigma_{H}(P):C(M;E)\otimes_{C(M)}C^{\ast}(Gr(H))\to C(M;F)\otimes_{C(M)}C^{\ast}(Gr(H)).$
The representation $\pi$ can be extended to a representation of $C^{\ast}(Gr(H)_{x})$ on $\mathcal{H_{\pi}}$, thus $\sigma_{H}(P)\otimes_{\pi}1$ on $(E_{x}\otimes C^{\ast}(Gr(H)_{x}))\otimes_{\pi}\mathcal{H_{\pi}}=E_{x}\otimes\mathcal{H_{\pi}}$ yields an operator
$\pi(\sigma_{H}(P,x)):E_{x}\otimes\mathcal{H}_{\pi}\to F_{x}\otimes\mathcal{H}_{\pi}.$
Remark 1.4.8.
Note that:
  • (i)
    The map from $\sigma_{H}(P)$ to $\pi(\sigma_{H}(P,x))$ is consistent with the $d\pi$ map defined in Definition 1.4.2 above by simple calculation, refer to [fischer2016quantization] Proposition 1.7.6.
  • (ii)
    The range and target map on the groupoid $Gr(H)$ denote the same map. Therefore, the groupoid $C^{\ast}$-algebra structure of $C^{\ast}(Gr(H))$ is a smooth family of group $C^{\ast}$-algebras $\{C^{\ast}{(Gr(H)_{x})}\}_{x\in M}$.
For left-invariant and homogeneous differential operators on a homogeneous Lie group, hypoellipticity is equivalent to the Rockland condition that we will define in the following theorem. The Rockland condition also aids in comprehending the definition of H-ellipticity.
Theorem 1.4.9.
[Rockland theorem, [melin1983parametrix]] The principal cosymbol $\sigma_{H}(P,x)$ of a differential operator $P$ at $x\in M$ is left invertible on $L^{2}(M)$ if and only if for every unitary non-trivial representation $\pi$ of the group $Gr(H)_{x}$, the (unbounded) operator $\pi(\sigma_{H}(P,x))$ is injective.
Definition 1.4.10.
An operator $P$ satisfying the condition in Theorem 1.4.9 is said to satisfy the Rockland condition.
Definition 1.4.11.
A differential operator $P$ of order $k$ on $M$ from vector bundle $E$ to $F$ is said to be H-elliptic if the principal cosymbol $\sigma_{H}^{k}(P,x)$ and $\sigma_{H}^{k}(P^{\ast},x)$ are both invertible at each point $x\in M$.
Remark 1.4.12.
  • According to the Rockland theorem, the condition in Definition 1.4.11 is equivalent to the condition that for all points $x\in M$, and for every non-trivial irreducible representation $\pi$ of $Gr(H)_{x}$, $\pi(\sigma_{H}(P,x))$ and $\pi(\sigma_{H}(P^{\ast},x))$ are injective.
  • In the pseudo-differential calculus, such a definition of H-ellipticity is equivalent to the invertibility of the cosymbol of the operator in the convolution distribution algebra $\mathcal{E}_{r}^{\prime}(Gr(H))/C^{\infty}_{p}(Gr(H);\Omega_{r})$, see Definition 1.6.6 and 1.6.7. For details, refer to [van2019groupoid].
1.5 Hilbert modules and $KK$-theory
In this section, we briefly introduce the Hilbert module theory and Kasparov’s $KK$-theory. See [blackadar1998k] for details.
Definition 1.5.1.
Let $B$ be a $C^{\ast}$-algebra. A pre-Hilbert module over $B$ is a right $B$-module $E$ equipped with a $B$-valued “inner product”, a function $\langle\cdot,\cdot\rangle:E\times E\rightarrow B$, with these properties:
  • (1)
[…]
  • (2)
    $\langle x,yb\rangle=\langle x,y\rangle b$ for all $x,y\in E,b\in B$.
  • (3)
    $\langle y,x\rangle=\langle x,y\rangle^{\ast}$.
  • (4)
    $\langle x,x\rangle\geq 0$, $\langle x,x\rangle=0$ if and only if $x=0$.
For $x\in E$, put $\lVert x\rVert=\lvert\langle x,x\rangle\rvert^{1/2}$. This is a norm on $E$. If $E$ is complete, $E$ is called a Hilbert module over $B$.
Definition 1.5.2.
Let $E$ be a Hilbert $B$-module. We define $\mathcal{L}(E)$ to be the set of all module homomorphisms $T:E\rightarrow E$ for which there is an adjoint module homomorphism $T^{\ast}:E\rightarrow E$ with $\langle Tx,y\rangle=\langle x,T^{\ast}y\rangle$ for all $x,y\in E$.
Proposition 1.5.3 ([blackadar1998k]).
Each operator in $\mathcal{L}(E)$ is bounded, and $\mathcal{L}(E)$ is a $C^{\ast}$-algebra with respect to the operator norm. In general, $\mathcal{L}(E)$ does not contain all bounded module-endomorphisms of $E$.
Definition 1.5.4.
  • If $x,y\in E$, let $\theta_{x,y}$ be the operator defined by $\theta_{x,y}(z)=x\langle y,z\rangle$ We have $\theta_{x,y}\in\mathcal{L}(E)$; in fact, $\theta_{x,y}^{\ast}=\theta_{y,x}$. If $T\in\mathcal{L}(E)$, then $T\theta_{x,y}=\theta_{Tx,y};\theta_{x,y}T=\theta_{x,T^{\ast}y}$, so the linear span of ${\theta_{x,y}}$ is an ideal in $\mathcal{L}(E)$. The space of compact operators $\mathcal{K}(E)$ is defined to be the closure of this linear span.
  • A regular operator on a Hilbert $B$-module $E$ is a densely defined operator $T$ on $E$ with densely defined adjoint $T^{\ast}$, such that $1+T^{\ast}T$ has dense range in $E$.
Remark 1.5.5.
If $T$ is a regular operator, then $(1+T^{\ast}T)^{-1}$ extends to an operator in $\mathcal{L}(E)$.
Definition 1.5.6.
Let $A$ and $B$ be graded $C^{\ast}$-algebras. $\mathbb{E}(A,B)$ is the set of all triples $(E,\phi,F)$, where $E$ is a countably generated graded Hilbert module over $B$, $\phi$ is a graded $\ast$-homomorphism from $A$ to $\mathcal{L}(E)$, and $F$ is an operator in $\mathcal{L}(E)$ of degree $1$, such that $[F,\phi(a)]$, $(F^{2}-1)\phi(a)$, and $(F-F^{\ast})\phi(a)$ are all in $\mathcal{K}(E)$ for all $a\in A$. The elements of $\mathbb{E}(A,B)$ are called Kasparov modules for $(A,B)$.
Definition 1.5.7.
  • Two triples $(E_{0},\phi_{0},F_{0})$ and $(E_{1},\phi_{1},F_{1})$ are unitarily equivalent if there is a unitary in $\mathcal{L}(E_{0},E_{1})$, of degree 00, intertwining the $\phi_{i}$ and $F_{i}$. Unitary equivalence is denoted $\sim_{u}$.
  • Let $IB$ denote the $C^{\ast}$-algebra of continuous functions from $[0,1]$ to $B$. A homotopy connecting $(E_{0},\phi_{0},F_{0})$ and $(E_{1},\phi_{1},F_{1})$ is an element $(E,\phi,F)$ of $\mathbb{E}(A,IB)$ for which $(E\hat{\otimes}_{f_{i}}B,f_{i}\circ\phi,{f_{i}}_{\ast}(F))\sim_{u}(E_{i},\phi_{i},F_{i})$, where $f_{i}$, for $i=0,1$, is the evaluation homomorphism from $IB$ to $B$. Homotopy equivalence is denoted $\sim_{h}$.
Definition 1.5.8.
The Kasparov $K$-group $KK(A,B)$ is defined to be the set of equivalence classes of $\mathbb{E}(A,B)$ under $\sim_{h}$.
Definition 1.5.9.
Let $A$ be a $C^{\ast}$-algebra.
  • A positive element $h\in A$ is strictly positive if $\phi(h)>0$ for every state $\phi:A\to\mathbb{C}$ of $A$. A $C^{\ast}$-algebra containing a strictly positive element will be called a $\sigma$-unital $C^{\ast}$-algebra.
  • A separable $C^{\ast}$-algebra is a $C^{\ast}$-algebra whose underlying topological space is a separable topological space.
Theorem 1.5.10 ([kasparov1988equivariant]).
Let $A$, $B$ be separable and $D$ be $\sigma$-unital graded $C^{\ast}$-algebras, let $(E_{1},\phi_{1},F_{1})$ be a Kasparov $(A,D)$-module and $(E_{2},\phi_{2},F_{2})$ be a Kasparov $(D,B)$-module, there is a Kasparov product for $(E_{1},\phi_{1},F_{1})$ and $(E_{2},\phi_{2},F_{2})$, which is a Kasparov $(A,B)$-module, unique up to homotopy $\sim_{h}$. The Kasparov product defines a bilinear function ${\otimes}_{D}:KK(A,D)\times KK(D,B)\rightarrow KK(A,B)$.
Corollary 1.5.11 ([kasparov1988equivariant]).
Let $A_{1}$, $A_{2}$, $B_{1}$, $B_{2}$, and $D$ be graded $C^{\ast}$-algebras with $A_{1}$ and $A_{2}$ separable and $B_{1}$ and $D$ $\sigma$-unital. There is a Kasparov product
$\otimes_{D}:KK(A_{1},B_{1}\hat{\otimes}D)\times KK(D\hat{\otimes}A_{2},B_{2})\to KK(A_{1}\hat{\otimes}A_{2},B_{1}\hat{\otimes}B_{2})$
which is is bilinear, contravariantly functorial in $A_{1}$ and $A_{2}$ and covariantly functorial in $B_{1}$ and $B_{2}$.
1.6 Heisenberg calculus for pseudo-differential operators
The normalization of the operators and inversion of cosymbol classes modulo smoothing terms are essential in our theory, but they are outside of the class of differential operators. In this section, we consider pseudo-differential operators.
General Lie groupoid
In [van2019groupoid], van Erp and Yuncken gave the pseudo-differential calculus on a filtered manifold through the Heisenberg tangent groupoid. We briefly review it here.
Definition 1.6.1.
Let $\mathcal{G}$ be a Lie groupoid. An $r$-fibered distribution on a Lie groupoid $\mathcal{G}$ is a continuous $C^{\infty}(\mathcal{G}^{(0)})$-linear map
$u:C^{\infty}(\mathcal{G})\to C^{\infty}(\mathcal{G}^{(0)}),$
where the $C^{\infty}(\mathcal{G}^{(0)})$-module structure on $C^{\infty}(\mathcal{G})$ is induced by the pullback of functions via $r$. The set of $r$-fibered distributions on $\mathcal{G}$ is denoted $\mathcal{E}^{\prime}_{r}(\mathcal{G})$. An $r$-fibered distribution $u$ determines a smooth family $\{u_{x}\}_{x\in\mathcal{G}^{(0)}}$ of distributions on the $r$-fibers $u_{x}\in\mathcal{E}^{\prime}(\mathcal{G}^{x})$,
$\langle u,f\rangle(x)=\langle u_{x},f|_{\mathcal{G}^{x}}\rangle,\quad f\in C^{\infty}(\mathcal{G}).$
Definition 1.6.2.
A subset $X$ of a groupoid $\mathcal{G}$ is proper if both $r:G\rightarrow M$ and $s:X\rightarrow M$ are proper maps. Denote by $\Omega_{r}$ and $\Omega_{s}$ the bundle of smooth densities tangent to the range and source fibers of $\mathcal{G}$ respectively. The space of smooth sections of $\Omega_{r}$ whose support is a proper subset of $\mathcal{G}$ is denoted $C_{p}^{\infty}(\mathcal{G};\Omega_{r})$.
Lemma 1.6.3 ([lescure2017convolution]).
$C_{p}^{\infty}(\mathcal{G};\Omega_{r})$ is a right ideal in $\mathcal{E}^{\prime}_{r}(\mathcal{G})$.
Definition 1.6.4.
$\mathbb{G}r(H)$ is defined to be the disjoint union of the osculating groupoid $Gr(H)$ with a family of pair groupoids indexed by ×=$\mathbb{R}^{\times}=\mathbb{R}\setminus\{0\}$,
$\mathbb{G}r(H)=Gr(H)\times\{0\}\sqcup(M\times M)\times\mathbb{R}^{\times},$
it is called the H-tangent groupoid of the osculating groupoid $Gr(H)$. The main result of [van2017tangent] is that there is a compatible global smooth structure on $\mathbb{G}r(H)$ that makes $\mathbb{G}r(H)$ into a Lie groupoid.
There is an action $\alpha:\mathbb{R}_{+}\to\text{Aut}(\mathbb{G}r(H))$, called the zoom action, defined to be a smooth one-parameter family of Lie groupoid automorphisms
$\alpha_{\lambda}(x,y,t)=(x,y,\lambda^{-1}t)\quad\text{if}$ $(x,y)\in M\times M,\ t\neq 0$
$\alpha_{\lambda}(x,\xi,0)=(x,\delta_{\lambda}(\xi),0)\quad\text{if}$ $(x,\xi)\in Gr(H),\ t=0$
where $\delta_{\lambda}$ is defined in Remark 1.1.4.
Definition 1.6.5.
A properly supported $r$-fibered distribution $\mathbb{P}\in\mathcal{E}^{\prime}_{r}(\mathbb{G}r(H))$ is called essentially homogeneous of weight $m\in\mathbb{R}$ if
${\alpha_{\lambda}}_{\ast}\mathbb{P}-\lambda^{m}\mathbb{P}\in C^{\infty}_{p}(\mathbb{G}r(H),\Omega_{r}).$
The space of essentially homogeneous distributions of weight $m$ is denoted as ${\Psi}_{H}^{m}(M)$.
Definition 1.6.6.
Define that:
  • A Schwartz kernel $P\in\mathcal{E}^{\prime}_{r}(M\times M)$ is an H-pseudo-differential kernel of order $\leq m$ if $P=\mathbb{P}|_{t=1}$ for certain $\mathbb{P}\in{\Psi}_{H}^{m}(M)$. The set of H-pseudo-differential kernels of order $\leq m$ will be denoted $\psi^{m}_{H}(M)$.
  • The cosymbol space is defined as:
    $\Sigma^{m}_{H}(M):=\{K\in\mathcal{E}^{\prime}_{r}(Gr(H))/C^{\infty}_{p}(Gr(H),\Omega_{r})|\ {\delta_{\lambda}}_{\ast}K=\lambda^{m}K\text{ for all }\lambda\in\mathbb{R}_{+}\}.$
  • The cosymbol map
    $\sigma_{m}:\psi^{m}_{H}(M)\to\Sigma^{m}_{H}(M)$
    is defined to select any $\mathbb{P}\in{\Psi}^{m}_{H}(M)$ with $\mathbb{P}|_{t=1}=P$ and set
    $\sigma_{m}(P):=\mathbb{P}|_{t=0}\text{ modulo }C_{p}^{\infty}(T_{H}M;\Omega_{r}).$
H-elliptic pseudo-differential operators can be characterized using the cosymbol defined above:
Definition 1.6.7.
An H-pseudo-differential operator $P\in\psi^{m}_{H}(M)$ is called H-elliptic if its principal cosymbol $\sigma_{m}(P)$ admits a convolution inverse in the cosymbol space $\Sigma^{m}_{H}(M)=\mathcal{E}^{\prime}_{r}(Gr(H))/$ $C_{p}^{\infty}(Gr(H);\Omega_{r})$.
Remark 1.6.8.
Note that:
  • (i)
    The Rockland condition is equivalent to the existence of a left parametrix for the cosymbol $\sigma_{H}(P)$. In the H-pseudo-differential calculus, if $P\in\psi_{H}^{m}(M)$ satisfies the Rockland criterion, then the operator $P$ itself acquires a left parametrix $Q\in\psi_{H}^{-m}(M)$ because of the existence of the left parametrix of the cosymbol $\sigma_{H}(P)$. Consequently, H-elliptic pseudo-differential operators are hypoelliptic.
  • (ii)
    As explained in Section 10 of [van2019groupoid], differential operators are special cases in pseudo-differential theory. General analysis in pseudo-differential calculus is crucial, although differential calculus facilitates an intuitive comprehension of the cosymbol and index theory of H-elliptic operators.
The cosymbol of a pseudo-differential operator can be viewed as a family of operators on the manifolds $Gr(H)_{x}$. So specifically, we need to analyze the case special case $\mathcal{G}=Gr(H)$.
Family of nilpotent Lie groups on a manifold
Note that $Gr(H)$ is a family of Lie groups over $M$ with a graded structure on $gr(H)$ and $\sigma_{H}(P)$ is a family of left-invariant pseudo-differential operators on $Gr(H)_{x}$.
So we conclude some results for left-invariant pseudo-differential operators on a nilpotent Lie groupoid with a graded vector bundle $\mathfrak{g}=\bigoplus_{i=1}^{n}\mathfrak{g}_{i}\to M$, such that $[\mathfrak{g}_{i},\mathfrak{g}_{j}]\subseteq\mathfrak{g}_{i+j}$ for all $i,j\in\mathbb{N}$ with $\mathfrak{g}_{k}=0$ if $k>n$. Assume that the bundle $\mathfrak{g}$ is equipped with an Euclidean metric such that different $\mathfrak{g}_{i}$’s are orthogonal to each other. The corresponding Lie group bundle is denoted by $\mathcal{G}$. The cosymbol operators are thus categorized as pseudo-differential operators in this calculus by considering $\mathcal{G}$ to be the osculating groupoid $Gr(H)$.
Definition 1.6.9.
Let $S^{m}(\mathcal{G})$ be the space of distributions $u\in\mathcal{E}^{\prime}(\mathcal{G})$ such that $u$ is transversal to the map $\pi:\mathcal{G}\to M$ and for each $\lambda\in\mathbb{R}^{+}$,
${\alpha_{\lambda}}_{\ast}u-\lambda^{m}u\in C_{c}^{\infty}(\mathcal{G}).$
Let $u\in S^{m}(\mathcal{G})$. For each $x\in M$, the operator can be defined as
$Op(u_{x}):C_{c}($ $\mathcal{G}_{x})$$\to$ $\ C_{c}(\mathcal{G}_{x})$
$f$ $\mapsto$ $(l\to(u_{x},f(l\cdot)))$
and they glue together to form an operator
$Op(u):C_{c}(\mathcal{G})\to C_{c}(\mathcal{G}).$
It can be extended to an operator on $C^{\ast}(\mathcal{G})$,
$\overline{Op(u)}:C^{\ast}(\mathcal{G})\to C^{\ast}(\mathcal{G}).$
Theorem 1.6.10 ([mohsen2020index]).
Let $u\in S^{m}(\mathcal{G})$, then
  • If $m\leq 0$, then $\overline{Op(u)}\in\mathcal{L}(C^{\ast}(\mathcal{G}))$;
  • If $m<0$, then $\overline{Op(u)}\in\mathcal{K}(C^{\ast}(\mathcal{G}))=C^{\ast}(\mathcal{G})$;
  • If $m>0$, and $u$ satisfy the Rockland condition, then $\overline{Op(u)}$ is a regular operator.
For left-invariant differential operators on $\mathcal{G}$, the Rockland condition facilitates the consideration of the $KK$-class:
Corollary 1.6.11 ([mohsen2020index]).
Let $P\in\mathcal{U}^{k}(\mathfrak{g})\otimes\text{Hom}(E,F)$ be a left-invariant differential operator on the bundle of Lie groups $\mathcal{G}$ satisfying the Rockland condition with $\alpha_{\lambda}(P)=\lambda^{k}(P)$ for $k\geq 0$. Then $\overline{P}$ is a regular operator in the sense of Baaj-Julg, and $\overline{P}^{\ast}=\overline{P^{\ast}}$. The operator
$D=\overline{P}/(1+\overline{P}^{\ast}\overline{P})^{1/2}\in\mathcal{L}(C(M;E)\otimes_{C(M)}C^{\ast}(\mathcal{G}),C(M;F)\otimes_{C(M)}C^{\ast}(\mathcal{G}))$
is a Fredholm operator in the sense of Kasparov $C^{\ast}(\mathcal{G})$-modules.
Corollary 1.6.12.
For any arbitrary H-elliptic operator $P$, the principal cosymbol $\sigma_{H}(P)$ represents a $KK$-element $[\sigma_{H}(P)]\in KK(\mathbb{C},C^{\ast}(Gr(H))\cong K_{0}(C^{\ast}(Gr(H)))$.
Proof.
By Corollary 1.6.11,
$\Big{[}C(M;E_{0}\oplus E_{1})\otimes_{C(M)}C^{\ast}(\mathcal{G}),\begin{pmatrix}0&\overline{P}^{\ast}/(1+\overline{P}\overline{P}^{\ast})^{1/2}\\ \overline{P}/(1+\overline{P}^{\ast}\overline{P})^{1/2}&0\end{pmatrix}\Big{]}$
represents an element in $KK(\mathbb{C},C^{\ast}(\mathcal{G}))\cong K_{0}(C^{\ast}(\mathcal{G}))$. Apply this result to $\sigma_{H}(P)$, the family of left invariant differential operators on $Gr(H)$, it represents an element in $KK(\mathbb{C},C^{\ast}(Gr(H))\cong K_{0}(C^{\ast}(Gr(H)))$. ∎
Later on we identify $KK(\mathbb{C},A)$ with $K_{0}(A)$ and $KK(A,\mathbb{C})$ with $K^{0}(A)$.
This framework can be extended to pseudo-differential operators [androulidakis2022pseudodifferential]. For a pseudo-differential operator $P$ on $M$, the principal cosymbol $\sigma_{H}(P)$ is an unbounded multiplier of the $C^{\ast}$-algebra $C^{\ast}(Gr(H))$, and the H-order of $P$ less or equal to 00 decide the boundedness or compactness of $P$ as follows:
Theorem 1.6.13 ([androulidakis2022pseudodifferential]).
Let $P$ be a pseudo-differential operators of H-order $k$ on $M$, then
  • If $k<0$, then $P$ extends to a compact operator $L^{2}(M,E)\to L^{2}(M,F)$;
  • If $k=0$, then $P$ extends to a bounded operator $L^{2}(M,E)\to L^{2}(M,F)$.
Remark 1.6.14.
In case $M$ is locally compact, denote the extension of $P$ to $L^{2}(M,E)$ by $\bar{P}$, then the result will becomes
  • If $k<0$, then $f\cdot\bar{P}$ is a compact operator for each $f\in C_{0}(M)$;
  • If $k=0$, then $f\cdot\bar{P}$ is a bounded operator for each $f\in C_{0}(M)$.
On a manifold $M$ with the filtration structure $H^{\bullet}$, a sequence of Sobolev spaces $\tilde{H}^{s}(M,E,F)$ are defined in [androulidakis2022pseudodifferential]. For a non-negative integer $s$, it encompasses the set of vector bundle $E$-valued distributions $u$ such that $Du\in L^{2}(F)$ for any differential operators $D\in\text{Diff}^{s}_{{H}}(M,E,F)$. It can be naturally extended to real numbers $s\geq 0$, and for $s<0$, $\tilde{H}^{s}(M)$ is defined to be the dual of $\tilde{H}^{-s}(M)$.
Theorem 1.6.15 ([androulidakis2022pseudodifferential]).
Let $M$ be a smooth compact manifold, and $P$ be a differential operator $C^{\infty}(M)\to C^{\infty}(M)$ of order $k$. Then the following are equivalent:
  • (1)
    For every $x\in M$ and $\pi\in gr(H)_{x}$, $\sigma(P,x,\pi)$ is injective on $C^{\infty}(H_{\pi})$.
  • (2)
    For every (or for some) $s\in\mathbb{R}$, and every distribution $u$ on $M$, $Pu\in\tilde{H}^{s}(M)$ implies $u\in\tilde{H}^{s+k}$.
  • (3)
    For every (or for some) $s\in\mathbb{R}$, the extension $\bar{P}:\tilde{H}^{s+k}(M)\to\tilde{H}^{s}(M)$ is left invertible modulo compact operators.
Corollary 1.6.16.
If $P$ is an H-elliptic differential operator of order $k$, $\bar{P}:\tilde{H}^{s}(M)\to\tilde{H}^{s-k}(M)$ is a Fredholm operator.
Definition 1.6.17.
Let $P$ be an H-elliptic differential operator of order $k$. For arbitrary $s\in\mathbb{R}$, $P$ can be extend to the Sobolev space $\tilde{H}^{s}(M)$,
$\bar{P}:\tilde{H}^{s}(M)\to\tilde{H}^{s-k}(M).$
We define the analytical index to be $\text{Ind}(P)=\dim\ker\bar{P}-\dim\ker\bar{P}^{\ast}$. Since H-elliptic operators are hypoelliptic, the kernel of $\bar{P}$ and $\bar{P}^{\ast}$ lie in $C^{\infty}(M)=\bigcap_{s\in\mathbb{R}}\tilde{H}^{s}(M)$, so such analytical index is well-defined.
Proposition 1.6.18 ([mohsen2022index]).
The analytical index is independent of the choice of $s$ and only depends on the symbol class $[\sigma_{H}(P)]\in K_{0}(C^{\ast}(Gr(H)))$.
Denote the map from the symbol class to the analytical index by
$\text{Ind}_{H^{\bullet}}:K_{0}(C^{\ast}(Gr(H)))\to\mathbb{Z}.$
Corollary 1.6.19.
In the $G$-invariant cases, all the $KK$-classes are replaced by equivariant $KK$-classes, and the analytical index $\text{Ind}(P)=[\text{Ker}\ P]-[\text{Coker}\ P]$ in the representation ring $R(G)$ only depends on the symbol class $[\sigma_{H}(P)]\in K^{G}_{0}(C^{\ast}(Gr(H)))$, i.e
$\text{Ind}_{H^{\bullet}}:K_{0}^{G}(C^{\ast}(Gr(H)))\to R(G).$
1.7 Groupoid approach and Connes-Thom’s isomorphism
Lemma 1.7.1.
Let $C$ be a family of $C^{\ast}$-algebras parametrized by $[0,1]$, such that the fiber over $(0,1]$ is the $C^{\ast}$-algebra $B$ and the fiber over 00 is the $C^{\ast}$-algebra $A$. There is an associated exact sequence
$0\to B\otimes C_{0}(0,1]\to C\xrightarrow{ev_{0}}A\to 0$
Then the induced map of $ev_{0}$ in $K$-theory is an isomorphism. Consequently, we can define a map in $K$-theory:
$\mu:=(ev_{1})_{\ast}\circ(ev_{0})_{\ast}^{-1}:K_{\ast}(A)\to K_{\ast}(B),$
where $ev_{t}$ is the restriction of $C$ to the $C^{\ast}$-algebra fibered over $\{t\}$.
Proof.
We have the $6$-term exact sequence in $K$-theory:
${K_{1}(B\otimes C_{0}(0,1])}$${K_{1}(C)}$${K_{1}(A)}$${K_{0}(A)}$${K_{0}(C)}$${K_{0}(B\otimes C_{0}(0,1]).}$
Since $C_{0}(0,1]$ is contractible, we see that $K(ev_{0}):K_{\ast}(C)\to K_{\ast}(A)$ is an isomorphism. ∎
Lemma 1.7.2.
If the global $C^{\ast}$-algebra $A$ can be decomposed as $A=\{A_{ij}\}_{(i,j)\in[0,1]\times[0,1]}$, a family of $C^{\ast}$-algebras over $[0,1]\times[0,1]$ with parameters $s,t\in[0,1]$, whose fibers are:
  • $A_{00}$: fibered over $\{0\}\times\{0\}$,
  • $A_{01}$: fibered over $\{0\}\times(0,1]$,
  • $A_{10}$: fibered over $(0,1]\times\{0\}$,
  • $A_{11}$: fibered over $(0,1]\times(0,1]$.
Then, we obtain a commutative diagram
${K_{\ast}(A_{00})}$${K_{\ast}(A_{10})}$${K_{\ast}(A_{01})}$${K_{\ast}(A_{11}).}$$\scriptstyle{\mu_{s=0}}$$\scriptstyle{\mu_{t=0}}$$\scriptstyle{\mu_{s=1}}$$\scriptstyle{\mu_{t=1}}$
in which we apply Lemma 1.7.1 to each side of $\partial{([0,1]\times[0,1])}$.
Proof.
Each component $A_{ij},i,j=0,1$ is a restriction of the global $C^{\ast}$-algebra $A$. The compositions along both side of the commutative diagram from $K_{\ast}(A_{00})$ to $K_{\ast}(A_{11})$ illustrate the induced map on $K$-theory, composition of the lift map from $A_{00}$ to the global algebra $A$, and the restriction map from $A$ to $A_{11}$. The result can then be derived as follows. ∎
Remark 1.7.3.
The same procedure can be performed by replacing $K_{\ast}(\cdot)$ with $KK_{\ast}(A^{\prime},\cdot)$ for $A^{\prime}$ an arbitrary $C^{\ast}$-algebra.
Theorem 1.7.4 (Connes-Thom’s isomorphism, [nistor2003index]).
Let $\mathcal{G}$ represent a bundle of simply-connected, solvable Lie group over $M$, then
$K_{i}(C^{\ast}(\mathcal{G}))\cong K_{i}(C_{0}(\mathfrak{g}^{\ast}))\cong K^{i}(\mathfrak{g}^{\ast}).$
Proposition 1.7.5.
There is a $KK$-element $[\mathcal{E}_{0}]\in KK(C^{\ast}(\mathcal{G}),C_{0}(\mathfrak{g}^{\ast}))$, called the Connes-Thom class, such that the $KK$-product with $[\mathcal{E}_{0}]$ over $C^{\ast}(\mathcal{G})$ yields the Connes-Thom’s isomorphism, i.e.
$K_{i}(C^{\ast}(\mathcal{G}))\xrightarrow[\cong]{\otimes_{C^{\ast}(G)}[\mathcal{E}_{0}]}K_{i}(C_{0}(\mathfrak{g}^{\ast})).$
Proof.
Referencing [fack1981connes], an analog of the standard Connes-Thom’s isomorphism for $C^{\ast}$-dynamic system $(A,\mathbb{R},\alpha)$:
$KK^{i}(B,A)\to KK^{i+1}(B,A\rtimes_{\alpha}\mathbb{R})$
can be represented in the Kasparov product with an element in $KK^{1}(A,A\rtimes_{\alpha}\mathbb{R})$. The isomorphism constructed in [nistor2003index] is a composite of such basic Connes-Thom’s isomorphisms. By associativity, the composition
$KK_{i}(A,C^{\ast}(\mathcal{G}))\cong KK_{i}(A,C_{0}(\mathfrak{g}^{\ast}))$
can be calculated by the Kasparov product with a certain element in $KK(C^{\ast}(\mathcal{G}),$ $C_{0}(\mathfrak{g}^{\ast}))$. ∎
Let us give a concrete construction of the Connes-Thom’s isomorphism in Theorem 1.7.4. Let $\mathcal{G}$ represent a bundle of nilpotent Lie groups over a compact manifold $M$. We can define a bundle of nilpotent Lie groups over $M\times[0,1]$, denoted by $\mathbb{G}$ and called the adiabatic groupoid for $\mathcal{G}$:
$\mathbb{G}=\mathcal{G}\times(0,1]\cup\mathfrak{g}\times\{0\},$
with its topology induced by the isomorphism
$\mathfrak{g}\times[0,1]$ $\to\mathbb{G}$
$(X,t)$ $\mapsto(\exp(tX),t)$ $\text{for}\ t>0,$
$(X,0)$ $\mapsto(X,0)$ $\text{for}\ t=0.$
Its $C^{\ast}$-algebra $C^{\ast}(\mathbb{G})$ lies in the exact sequence
$0\to C_{0}(0,1]\otimes C^{\ast}\mathcal{G}\to C^{\ast}(\mathbb{G})\xrightarrow{ev_{0}}C_{0}(\mathfrak{g}^{\ast})\to 0$
By the arguments in Lemma 1.7.1, the map $ev_{0}$ induces an isomorphism in Kasparov’s $KK$-theory.
Lemma 1.7.6 ([nistor2003index], Theorem 2.22).
For arbitrary $C^{\ast}$-algebra $A$, the composition
$(ev_{1})_{\ast}\circ(ev_{0})_{\ast}^{-1}:KK(A,C_{0}(\mathfrak{g}^{\ast}))\to KK(A,C^{\ast}(\mathbb{G}))\to KK(A,C^{\ast}\mathcal{G})$
is an isomorphism. It is consistent with the inverse of the Connes-Thom’s isomorphism and equal to the $KK$-product with $[ev_{0}]^{-1}\in KK(C_{0}(\mathfrak{g}^{\ast}),C^{\ast}(\mathbb{G}))$ and $[ev_{1}]\in KK(C^{\ast}(\mathbb{G}),C^{\ast}(\mathcal{G}))$.
For each $x\in M$, The Lie group $Gr(H)_{x}$ is nilpotent and thus solvable; it is also simply-connected due to its construction. Thus, let $\mathcal{G}=Gr(H)_{x}$, we have
Corollary 1.7.7.
For any $C^{\ast}$-algebra $A$, there is an isomorphism
$KK(A,C_{0}(gr(H)^{\ast}))\cong KK(A,C^{\ast}(Gr(H))).$
It is also called the Connes-Thom’s isomorphism. We will use this isomorphism in Chapter 3 and 4 in the H-elliptic theory and the transversally H-elliptic theory.
Chapter 2 Elliptic operators and transversally elliptic operators
2.1 Elliptic operators from $KK$-theoretic viewpoint
This section introduced the index theory of elliptic operators discussed in [kasparov2017elliptic], elucidating the symbol class and index class in $KK$-homology.
Definition 2.1.1 ([kasparov2017elliptic]).
A pseudo-differential operator $P$ of order 00 on $M$ is said to be elliptic if
$\lVert\sigma(P,x,\xi)\sigma(P^{\ast},x,\xi)-1\rVert\to 0,$
$\lVert\sigma(P^{\ast},x,\xi)\sigma(P,x,\xi)-1\rVert\to 0$
when $\xi\to\infty$.
Remark 2.1.2.
The definition asserting that $\sigma(P)$ is invertible at infinity aligns well with this definition. If $\sigma(P)$ is invertibility at infinity in $\xi$ , it can be normalized by replacing it with $((\sigma(P)\sigma(P^{\ast}))^{-1/2}\sigma(P)$. The new symbol is unitary at infinity in $\xi$ and homotopic to the initial one. Moreover, upon replacing $P$ with the self-adjoint operator
$\begin{pmatrix}0&P^{\ast}\\ P&0\end{pmatrix},$
any elliptic operator $P$ can be reformulated into a zero-order elliptic self-adjoint operator on $\mathbb{Z}_{2}$-graded bundle $E=E_{0}\oplus E_{1}$, which we still denote by $P$, the definition of ellipticity can be replaced as
$\lVert\sigma(P,x,\xi)^{2}-1\rVert\to 0.$ (*)
Henceforth, we assume that elliptic operators satisfy the estimate (*2.1.2). This definition helps to give a direct comprehension of the $KK$-classes related to elliptic operators.
Definition 2.1.3.
We define that:
  • The symbol class for a zero-order elliptic operator is defined by the $KK$-pairing
    $[\sigma(P)]=[(C_{0}(T^{\ast}M;p^{\ast}(E)),\sigma(P)]\in K_{0}^{G}(C_{0}(T^{\ast}M)),$
    because $1-\sigma^{2}(P)\in C_{0}(T^{\ast}M;p^{\ast}(E))$ is a compact multiplier on $C_{0}(T^{\ast}M;p^{\ast}(E))$;
  • The index class is defined by the $KK$-pairing
    $[P]=[(L^{2}(E),P)]\in K_{G}^{0}(\mathbb{C}).$
The index theorem is expressed as:
Theorem 2.1.4 ([kasparov2022k]).
Let $M$ be a compact Riemannian manifold and $G$ be a Lie group acting isometrically on $M$. Let $P$ be a bounded elliptic operator of order 00 on $M$, then
$[P]=[\sigma(P)]\otimes_{C_{0}(T^{\ast}M)}[\mathcal{D}_{M}]\in KK_{G}(\mathbb{C},\mathbb{C}),$
where $[\mathcal{D}_{M}]\in K^{0}_{G}(C_{0}(TM))$ denotes the $KK$-class of the Dolbeault operator, the $KK$-product on the right-hand side is considered that
$KK_{G}(\mathbb{C},C_{0}(T^{\ast}M))\times KK_{G}(C_{0}(T^{\ast}M),\mathbb{C})\xrightarrow{{\otimes}_{C_{0}(T^{\ast}M)}}KK_{G}(\mathbb{C},\mathbb{C}).$
Here, the Dolbeault operator represents the canonical Dirac operator $D$ on $TM$ with the symbol
$\sigma_{D}(x,\xi,\chi,\zeta)=(\text{ext}\ \zeta+\text{int}\ \zeta)+i(\text{ext}\ \chi-\text{int}\ \chi)).$
The coordinates in the cotangent space $T^{\ast}_{(x,\xi)}(TM)$ is denoted as $(\chi,\zeta)$.
Proof.
The proof is given in [kasparov2022k], and we briefly conclude the proof here.
Primarily, the proof constructs a rotation homotopy between the symbol of two operators, representing the $KK$-classes of the symbol class and index class.
More explicitly, considering the Kasparov product of the Bott–Dirac element $[\mathcal{C}]$ and $[P]$ yields an operator that consolidates $\mathcal{C}\hat{\otimes}1$ and the lifting $\tilde{P}$ of the operator $P$ to $TM$, as $\tilde{P}$ denotes a $KK$-theoretic connection of $P$. The resulting operator acts on $TM$.
The Kasparov product of $\sigma_{F}$ and $[\mathcal{D}_{M}]$ can be represented as the operator combining the lifting $\widetilde{\mathcal{D}_{M}}$ and $1\hat{\otimes}\sigma_{F}$, functioning on $T^{\ast}M$.
Through a rotation of $(\xi,\chi)$-coordinate, the symbols of these operators are rendered homotopic. Given that the Bott–Dirac element $[\mathcal{C}]$ represent the trivial element in $KK$-homology, we see that
$[P]=[\sigma(P)]\otimes_{C_{0}(T^{\ast}M)}[\mathcal{D}_{M}].$
2.2 Transversally elliptic operators from $KK$-theoretic viewpoint
For transversally elliptic operators defined in Definition 2.2.2, it turns out that the direct replacement of the $C_{0}(TM)$ space by $C_{0}(T_{G}M)$ space is not enough. As discussed in [berline1996chern], additional conditions are necessary to identify the “good” symbols due to the inadequate nature of the algebra $C_{0}(T_{G}M)$ for formulating an index theorem.
Kasparov’s index theorem for transversally elliptic operators in $KK$-theory is introduced in [kasparov2022k], presenting his transversally elliptic theory, including the concept of the tangent Clifford symbol, which underpins the definition of transversal H-ellipticity.
Let $M$ be a compact Riemannian manifold equipped with a smooth, isometric action of a Lie group $G$. Let $x\in M$ and $f_{x}:G\to M$ illustrate the map defined by $g\mapsto g(x)$. The Lie algebra of the group $G$ is denoted as $\mathfrak{g}$, and accordingly, $f^{\prime}_{x}:\mathfrak{g}\to T_{x}M$, which accounts for the derivative of $f_{x}$ at $e$. The $G$-action on $M$ is proper, and thus, there exists a $G$-invariant Riemannian metric on $\mathfrak{g}_{M}=M\times\mathfrak{g}$. There is a $G$-invariant quadratic form $q$ on covectors $\xi\in T_{x}^{\ast}M$ by
$q_{x}(\xi)=\lVert{f^{\prime}_{x}}^{\ast}(\xi)\rVert^{2},$
and a $G$-invariant map
$\varphi_{x}:=f^{\prime}_{x}{f^{\prime}_{x}}^{\ast}:T_{x}M\to T_{x}M.$
Remark 2.2.1.
Note that $\xi\in T_{x}M$ is orthogonal to the $G$-orbit passing through $x$ if and only if $q_{x}(\xi)=0$, or if and only if $\varphi_{x}(\xi)=0$, denote the space of such $\xi$ by $T_{G}M$, it is called the transversal part of $TM$ with $G$-action.
Kasparov’s definition of transversally elliptic operators is stated as follows:
Definition 2.2.2.
A $G$-invariant pseudo-differential operator $P$ of order 00 on a vector bundle $E$ over $M$ is called transversally elliptic if its symbol $\sigma(P)$ satisfies the following condition: for any $\epsilon\geq 0$, there exists $c>0$ such that for any $x\in M$, $\xi\in T^{\ast}_{x}M$,
$\lVert\sigma(P)^{2}(x,\xi)-1\rVert\leq c(1+q_{x}(\xi))(1+\lVert\xi\rVert^{2})^{-1}+\epsilon.$
Remark 2.2.3.
  • For a $G$-invariant properly supported normalized pseudo-differential operator $P$ of order 00, the original definition of the transversally elliptic is that if $\lVert\sigma(P,x,\xi)^{2}-1\rVert\to 0$ uniformly in $x\in M$ as $(x,\xi)\to\infty$ in $T_{G}M$. Note that we have replaced the operator $P$ by the normalization
    $D=\begin{pmatrix}0&P^{\ast}\\ P&0\end{pmatrix},$
    such that it acts on the $\mathbb{Z}_{2}$-graded bundle $E=E_{0}\oplus E_{1}$ of 00-order and self-adjoint. By [kasparov2017elliptic], Lemma 6.2, these two definitions coincide for 00-order pseudo-differential operators.
  • In Definition 7.1 of [kasparov2022k], there is a space of negative order symbol $\mathcal{S}_{tr}(M)$, defined as the sub-algebra of $C_{b}(TM)$ comprising functions $b(x,\xi)$ with the pseudo-differential estimate, and such that for any $\epsilon>0$,
    $|b(x,\xi)|\leq c(1+q_{x}(\xi))(1+\lVert\xi\rVert^{2})^{-1}+{\epsilon}$
    for some certain $c>0$. The symbol $\sigma(P)$ of a transversally elliptic operator will induce a class $[\sigma(P)]\in KK_{G}(\mathbb{C},\mathcal{S}_{tr}(M))$.
We need to consider the tangent-Clifford symbol of $P$, which is vital for the transversally H-elliptic analysis.
Let $V$ be a real vector bundle over $M$ equipped with a $G$-invariant Riemannian metric. Take $\text{Cliff}(V,Q)$ the Clifford algebra bundle associated with the quadratic form $Q(v)=\lVert v\rVert$ on $V$. The norm on fibers of the Clifford bundle $\text{Cliff}(V,Q)$ is defined by its embedding into the bundle of endomorphisms $\mathcal{L}(\Lambda^{\ast}(V))$ of the exterior algebra of $V$.
Definition 2.2.4.
  • Define $Cl_{V}(M)$ to be the complexification of the algebra of continuous sections of $\text{Cliff}(V,Q)$ over $M$, forming a $C^{\ast}$-algebra with the sup-norm on sections.
  • For the manifold $M$ with group $G$-action, define the bundle
    $\Gamma=\{f^{\prime}(s)|\ s\in C_{0}(M;M\times\mathfrak{g})\},$
    where $f^{\prime}$ denotes the collection of maps $f^{\prime}_{x}:\mathfrak{g}\to T_{x}M$ defined at beginning of this section. Then $Cl_{\Gamma}(M)$ is the algebra of continuous sections of the sub-algebra of $\text{Cliff}(TM)$ which is generated by vectors tangent to the orbit direction.
Definition 2.2.5.
Let $P$ be a 00-order $G$-invariant transversally elliptic operator on $E$ over $M$.
  1. (a)
    The index class $[P]$ is defined as $[(L^{2}(E),P)]\in K^{\ast}(C^{\ast}(G))$.
  2. (b)
    The tangent-Clifford symbol class
    $[\sigma^{tcl}(P)]:=[(C_{0}(T^{\ast}M;p^{\ast}E)\hat{\otimes}_{C(M)}Cl_{\Gamma}(M),\sigma^{tcl}(P))]\in KK_{G}(\mathbb{C},Cl_{\Gamma}(TM)),$
    is defined by the operator
    $\sigma^{tcl}(P)(x,\xi):=\sigma(P)(x,\xi)\hat{\otimes}1+(1-\sigma(P)^{2})^{1/2}\hat{\otimes}f_{\Gamma}(x,\xi)\in\mathcal{L}((C_{0}(T^{\ast}M;p^{\ast}E)\hat{\otimes}_{C(M)}Cl_{\Gamma}(M)),$
    where
    • $Cl_{\Gamma}(TM):=Cl_{\Gamma}(M)\otimes_{C(M)}C_{0}(TM)$,
    • $f_{\Gamma}(x,\xi)=\varphi_{x}(\xi)/(1+q_{x}(\xi))^{1/2}$ denotes a covector field on $T^{\ast}M$, which represents the Clifford multiplication operator.
    The tangent-Clifford symbol class is denoted by $[\sigma^{tcl}(P)]\in KK_{G}(\mathbb{C},Cl_{\Gamma}(TM))$.
Remark 2.2.6.
By the $KK^{G}$-equivalence in some sense between the algebras $Cl_{\Gamma}(TM)$ and $\mathcal{S}_{tr}(M)$([kasparov2022k]), a one-to-one correspondence exists between the symbol class $[\sigma(P)]\in KK_{G}(\mathbb{C},\mathcal{S}_{tr}(M))$ and the tangent-Clifford symbol class, $[\sigma^{tcl}(P)]\in KK_{G}(\mathbb{C},Cl_{\Gamma}(TM))$.
The definition of $\mathcal{S}_{tr}(M)$ space depends on the inequality estimate, rendering it unsuitable for the transversally H-elliptic case as $C^{\ast}(Gr(H))$ is noncommutative. However, the tangent-Clifford symbol class only depends on the orbit direction in $T_{x}M$ for each $x\in M$, which is considerable in the transversally H-elliptic case.
Based on the $KK^{G}$-equivalence, we can always analyze the transversally elliptic symbols by using the tangent-Clifford symbol class $[\sigma^{tcl}(P)]$ in the following contents.
Definition 2.2.7.
The Clifford Dolbeault element $[D^{cl}_{M,\Gamma}]\in K^{0}(C^{\ast}(G,Cl_{\Gamma}(TM)))$ is defined to be the $KK$-element $[(\mathcal{H},\pi,F)]$, where
  • $\mathcal{H}=L^{2}(\Lambda_{\mathbb{C}}^{\ast}(TM))$;
  • $F=\mathcal{D}_{M}$ is the same operator as the Dolbeault element $[\mathcal{D}_{M}]\in K^{0}_{G}(C_{0}(TM))$;
  • The representation $\pi$ is extended from $C_{0}(TM)$ to $C^{\ast}(G,Cl_{\Gamma}(TM))$, by $\zeta\to(-\text{ext}(\zeta)+\text{int}(\zeta))$ for $\zeta$ a section of $\Gamma$, and the natural group action $G$ on $L^{2}(\Lambda_{\mathbb{C}}^{\ast}(TM))$.
From these definitions, the transversally elliptic index theorem is stated as follows:
Theorem 2.2.8 ([kasparov2022k]).
Let $M$ be a compact Riemannian manifold and $G$ be a Lie group which acts on $M$ isometrically, Let $P$ represent a $G$-invariant, $L^{2}$-bounded transversally elliptic operator on $M$ of order 00, then
$[P]=j^{G}([\sigma^{tcl}(P)])\otimes_{C^{\ast}(G,Cl_{\Gamma}(TM))}[\mathcal{D}_{M,\Gamma}^{cl}],$
where $j^{G}$ denote the natural map from $G$-equivariant $KK$-homology to $KK$-homology of crossed product with $G$, refer to [blackadar1998k]. The $KK$-product on the right-hand side is considered
$KK(C^{\ast}(G),C^{\ast}(G,Cl_{\Gamma}(TM))\times KK(C^{\ast}(G,Cl_{\Gamma}(TM),\mathbb{C})\xrightarrow{\otimes_{C^{\ast}(G,Cl_{\Gamma}(TM)}}KK(C^{\ast}(G),\mathbb{C}).$
Chapter 3 H-elliptic index theory
In this chapter, we introduce the H-elliptic index theorem. In [mohsen2022index], Mohsen constructed a diagram of groupoids over $[0,1]^{2}$, such that the groupoid structure on each side independently forms a Lie groupoid structure. A $C^{\ast}$-algebra is defined on this groupoid, analyzed and integrated to deduce the index theorem for H-elliptic differential operators.
3.1 The groupoid approach for H-elliptic operators
Let $M$ be a smooth closed filtered manifold. We consider the global groupoid structure which constructed over $[0,1]^{2}$ with units in $M\times[0,1]^{2}$ as follows:
  • the pair groupoid $M\times M$ if $(t,s)\in(0,1]^{2}$;
  • the tangent bundle $TM$ if $(t,s)\in(0,1]\times\{0\}$;
  • the bundle of osculating groups $Gr(H):=\bigsqcup_{x\in M}Gr(H)_{x}$ if $(t,s)\in\{0\}\times(0,1]$;
  • the bundle of abelian groups $gr(H):=\bigsqcup_{x\in M}gr(H)_{x}$ if $(t,s)=(0,0)$.
For some fixed $t_{0}$, the tangent bundle $TM$ over $(t_{0},0)$ and pair groupoid $M\times M$ over $(t_{0},s)$ for $s\in(0,1]$ can be attached by the convergence relation
$x_{s},y_{s}\in M\times M,\ V\in T_{x}M,\ (x_{s},y_{s},s)\to(V,0)\ \ \text{if}\ \ s^{-1}(y_{s}-x_{s})\to V$
to form the tangent groupoid
$\mathbb{T}M=TM\times\{0\}\cup M\times M\times(0,1].$
By the same way, the groupoids constructed above can be attached by the diagram
${gr(H)}$${TM}$${Gr(H)}$${M\times M,}$$\scriptstyle{s^{-1}}$$\scriptstyle{\delta_{t}^{-1}}$$\scriptstyle{s^{-1}}$$\scriptstyle{\delta_{t}^{-1}}$
where $s^{-1}$ and $\delta_{t}^{-1}$ indicates the the convergences relations.
Each side of the diagram independently constitutes a groupoid structure. At $t=1$, the attached groupoid is the tangent groupoid of $TM$, and at $t=0$, it is the adiabatic groupoid constructed in Section 1.7. At $s=0$, it is perceived as the $C^{\infty}(M\times\mathbb{R})$-submodule of $C^{\infty}(M\times\mathbb{R};TM\times\mathbb{R})$ consisting of sections:
$\big{\{}\mathbb{X}\in C^{\infty}(M\times\mathbb{R};TM\times\mathbb{R})\big{|}\ \partial_{t}^{k}\mathbb{X}|_{t=0}\in C^{\infty}(M;H^{k})\text{ for all }t\geq 0\big{\}}.$
At $s=1$, the attached groupoid is the Heisenberg tangent groupoid, or is called the holonomy groupoid of the adiabatic foliation $\mathfrak{a}H$ associated with $H$. Further information on holonomy groupoid and adiabatic foliation is available in [androulidakis2022pseudodifferential]
We apply the restriction of the global groupoid $C^{\ast}$-algebra to each side. Then, as discussed in Section 1.7, for a family of $C^{\ast}$-algebras over $[0,1]$, with $B$ fibered over $(0,1]$ and $A$ over 00, we obtain an exact sequence that induces a map in $K$-theory or $KK$-theory that
$(ev_{1})_{\ast}\circ(ev_{0})_{\ast}^{-1}:K_{\ast}(A)\to K_{\ast}(B).$
Using this construction for the $C^{\ast}$-algebra of the groupoids over each side, by Lemma 1.7.2, we can obtain a commutative diagram:
${K^{0}(gr(H)^{\ast})}$${K^{0}(T^{\ast}M)}$${K_{0}(C^{\ast}(Gr(H)))}$${\mathbb{Z}.}$$\scriptstyle{\mu_{t=0}}$$\scriptstyle{\mu_{s=0}}$$\scriptstyle{\mu_{t=1}}$$\scriptstyle{\mu_{s=1}}$
Remark 3.1.1.
Note that the $\mu_{s=0}$ is induced by the attaching between $gr(H)^{\ast}$ and $T^{\ast}M$ facilitated by $\delta_{t}^{-1}$. Subsequently, we explore the map between the $KK$-groups of $C^{\ast}$-algebras induced by the isomorphism from the bundle $TM$ to $gr(H)$ via orthogonal decomposition—such an isomorphism is called a splitting isomorphism ([van2017tangent]). The induced map in $K$-theory remains invariant regardless of the selected isomorphism, leading to a concurrence of these two induced maps through direct computation.
Lemma 3.1.2.
The $\mu_{t=0}$ map is the Connes-Thom’s isomorphism in Corollary 1.7.7. The $\mu_{t=1}$ map is the ordinary analytical index map given by Connes in [connes1994noncommutative]. The map $\mu_{s=1}$ is the analytical index map $\text{Ind}_{H^{\bullet}}$ in the Heisenberg calculus, which is defined in Section 1.6.
Proof.
By the concrete construction for Connes-Thom’s class before Lemma 1.7.6, the $\mu_{t=0}$ map is the Connes-Thom’s isomorphism. The $\mu_{t=1}$ is same as the Atiyah-Singer analytical index map $\text{Ind}_{a}$ defined in Lemma 2.5.6 in [connes1994noncommutative]. The last part is proved in Section 3 in [van2010atiyah1]. ∎
The commutativity of the diagram and Lemma 4.1 in [mohsen2022index] implies that
Theorem 3.1.3 ([mohsen2022index]).
The index map $\text{Ind}_{H^{\bullet}}:K_{0}(Gr(H))\to\mathbb{Z}$ can be evaluated as
$\text{Ind}_{H^{\bullet}}=\text{Ind}_{a}\circ\mu_{s=0}\circ\mu_{t=0}^{-1},$
where $\mu_{t=0}^{-1}:K_{0}(C^{\ast}(Gr(H)))\to K^{0}(gr(H)^{\ast})$ representing Connes-Thom’s isomorphism. This constitutes the index theorem for H-elliptic differential operators.
An alternative formulation of this index theorem will be given in the next section and be proved in Chapter 5 using the $KK$-theoretic argument.
3.2 Index theorem for H-elliptic operators from $KK$-theoretic viewpoint
This section derives the index theorem for H-elliptic differential operators from the $KK$-theoretic perspective, expressed as a $KK$-product representation. We focus on the index of operators in $KK_{0}$, the $KK_{1}$ index for self-adjoint operators on non-graded bundles can be analogously analyzed. The H-elliptic operators can be analyzed using Kasparov’s $KK$-theory.
In the $G$-invariant cases, the Lie group $G$ acts on $TM$ induced by the $G$-action on $M$. We assume that $G$ preserves the H-filtration structure. Then, we have
Lemma 3.2.1.
There is a natural $G$-action on the bundle $gr(H)$ by $g[v]=[gv]$ for arbitrary $g\in G$ and $v\in H^{i}_{x}$. It induces a $G$-action on the bundle of groups $Gr(H)$ by the exponential map.
For a general H-elliptic differential operator $P_{0}$, we consider the bounded transform
$P=\begin{pmatrix}0&P_{0}^{\ast}/(1+P_{0}P_{0}^{\ast})\\ P_{0}/(1+P_{0}^{\ast}P_{0})&0\end{pmatrix},$
which is a $G$-invariant Fredholm operator of order 00. $[a,P]$ and $a(1-P^{2})$ are operators with negative order symbols, so they are both compact. Hence, we can define that
Definition 3.2.2.
The index class $[P_{0}]$ of $P_{0}$ is defined to be the $KK$-class $[(L^{2}(E_{0}\oplus E_{1}),P)]\in KK_{G}(\mathbb{C},\mathbb{C})\cong R(G)$.
Remark 3.2.3.
Henceforth, we will always assume that the considered operator $P_{0}$ is transformed as above to a self-adjoint operator $P$ of 00-order. We identify the index classes $[P]=[P_{0}]$.
By Proposition 1.6.18 , a 00-order H-elliptic operator $P$ induces the symbol class
$[\sigma_{H}(P)]\in KK(\mathbb{C},C^{\ast}(Gr(H)))\cong K_{0}(C^{\ast}(Gr(H))).$
and in the $G$-invariant case,
$[\sigma_{H}(P)]\in KK^{G}(\mathbb{C},C^{\ast}(Gr(H)))\cong K^{G}_{0}(C^{\ast}(Gr(H))).$
can be defined in a similar manner.
Apply $\mathcal{G}=Gr(H)$ in Proposition 1.7.5, the Connes-Thom’s isomorphism can be represented by a $KK$-product with a certain element $[\mathcal{E}_{0}]\in KK(C^{\ast}(Gr(H)),$ $C_{0}(gr(H)^{\ast}))$. Let
$[\mathcal{E}]=[\mathcal{E}_{0}]\otimes_{C^{\ast}(gr(H)^{\ast})}[ev_{0}]^{-1}\otimes[ev_{1}]\in KK(C^{\ast}(Gr(H)),C_{0}(T^{\ast}M)),$
where $[ev_{1}]$ and $[ev_{0}]$ represent the $KK$-element given by the $C^{\ast}$-algebra homomorphisms $ev_{1}$ and $ev_{0}$ in $s=0$ side.
Theorem 3.2.4.
The analytical index map $\text{Ind}_{H^{\bullet}}$ for H-elliptic operators can be expressed by $KK$-products as
$\text{Ind}_{H^{\bullet}}=\otimes_{C^{\ast}(Gr(H))}[\mathcal{E}]\otimes_{C_{0}(T^{\ast}M)}[\mathcal{D}_{M}].$
Proof.
The Kasparov product with $[\mathcal{E}]$ facilitates an isomorphism between $KK(\mathbb{C},C^{\ast}(Gr(H))$ and $KK(\mathbb{C},C_{0}(T^{\ast}M))$, which represents the composition of the Connes-Thom’s isomorphism and the induced isomorphism of $K^{0}$ between $T^{\ast}M$ and $gr(H)^{\ast}$, i.e.,
$\mu_{s=0}\circ\mu^{-1}_{t=0}=\otimes_{C^{\ast}(Gr(H))}[\mathcal{E}].$
The Kasparov product with $[\mathcal{D}_{M}]$ gives the ordinary analytical index map, i.e.,
$\text{Ind}_{a}=\mu_{t=1}=\otimes_{C_{0}(T^{\ast}M)}[\mathcal{D}_{M}].$
Apply these two equations to Theorem 3.1.3, we see that
$\text{Ind}_{H^{\bullet}}$ $=\text{Ind}_{a}\circ\mu_{s=0}\circ\mu_{t=0}^{-1}$
$=\otimes_{C^{\ast}(Gr(H))}[\mathcal{E}]\otimes_{C_{0}(T^{\ast}M)}[\mathcal{D}_{M}].$
By the natural construction of $G$-action on $gr(H)$, the Connes-Thom’s isomorphism and the induced isomorphism $\mu_{s=0}$ from $gr(H)$ to $TM$ is $G$-invariant. Consequently,
Proposition 3.2.5.
$[\mathcal{E}]$ represent a $G$-invariant $KK$-class in $KK_{G}(C^{\ast}(Gr(H)),C_{0}(T^{\ast}M))$. The Kasparov product with $[\mathcal{E}]$ provides an isomorphism between $KK_{G}(\mathbb{C},$ $C^{\ast}(Gr(H))$ and $KK_{G}(\mathbb{C},C_{0}(T^{\ast}M))$.
Thus, the index theorem for H-elliptic operators in $KK$-theory can be expressed as
Theorem 3.2.6.
Let $M$ be a compact filtered Riemannian manifold and $G$ be a Lie group, which isometrically acts on $M$ and preserves the filtered structure. If $P_{0}$ be an H-elliptic $G$-invariant differential operator on $M$, let
$P=\begin{pmatrix}0&P_{0}^{\ast}/(1+P_{0}P_{0}^{\ast})\\ P_{0}/(1+P_{0}^{\ast}P_{0})&0\end{pmatrix}$
be the bounded transform of $P_{0}$ of order 00, then we have
$[P]=[\sigma_{H}(P)]\otimes_{C^{\ast}(Gr(H))}[\mathcal{E}]\otimes_{C_{0}(T^{\ast}M)}[\mathcal{D}_{M}],$
where $[\mathcal{D}_{M}]\in K^{0}_{G}(C_{0}(TM))$ denote the $KK$-class of the Dolbeault operator.
The $KK$-product on the right-hand side is executed over
$KK_{G}(\mathbb{C},C^{\ast}(Gr(H))$ $\otimes_{C^{\ast}(Gr(H))}KK_{G}(C^{\ast}(Gr(H)),C_{0}(T^{\ast}M))$
$\otimes_{C_{0}(T^{\ast}M)}KK_{G}(C_{0}(T^{\ast}M),\mathbb{C}).$
The proof will be given in chapter 5.
Later, we will use the following lemma in the proof of Theorem 3.2.6 and the main Theorem 5.2.2.
Lemma 3.2.7.
Let $P$ be the bounded transform of $P_{0}$ as in Theorem 3.2.6. For the multiplier $\sigma_{H}(P)$ of $C^{\ast}(Gr(H))$, representing an element of $KK_{G}(\mathbb{C},C^{\ast}(Gr(H)))$, the product with $[\mathcal{E}]$ induces the transformation as the following three steps:
  • (i)
    Freezing the coefficient of the multiplier $\sigma_{H}(P)(x)=P_{x}$ to an element in $\mathcal{U}(gr(H)_{x})$ for each $x\in M$, which represents a differential operator on $C^{\ast}(gr(H)_{x})$ as $gr(H)_{x}$ is a vector space.
  • (ii)
    Choose a splitting isomorphism between the bundle $gr(H)$ and $TM$ (herein, we presume it to be the orthonormal decomposition in Definition 4.1.1), the differential operator family on $C^{\ast}(gr(H)_{x})$ can be transformed into a family of operators $P^{\prime}=\{P^{\prime}_{x}\}_{x\in M}$ acting on $C^{\ast}(T_{x}M)$ for each $x\in M$.
  • (iii)
    Take the Fourier transform from $C^{\ast}(TM)$ to $C_{0}(T^{\ast}M)$, the family of operators $P^{\prime}$ will be transformed into the multiplication operator $\sigma^{\prime}(P)$ on $C_{0}(T^{\ast}M)$ (If $P$ act on bundle $E$, we need to tensor product endomorphisms on $E$), which represent an element in $K^{0}(T^{\ast}M)$
This procedure yields a multiplicative operator $\sigma^{\prime}(P)(x,\xi)$ on $C_{0}(TM)$, which is essentially the Fourier transform of the freezing coefficient operators of the H-principal component of the original operator $P$.
Proof.
First, let $P$ be the H-elliptic $G$-invariant differential operator $P_{0}$.
The product with $[\mathcal{E}]$ is the composition of isomorphisms $\mu_{t=0}^{-1}$ and $\mu_{s=0}$. The first one is induced from the adiabatic groupoid for $Gr(H)_{x}$ as in Proposition 1.7.5. Let $P_{0,x}$ be the freezing coefficient of $P_{x}$ to $e\in Gr(H)_{x}$, by the arguments in [van2015groupoid], we see that on this adiabatic groupoid, the family $\mathbb{P}_{t}$ defined by
$\mathbb{P}_{t}=\begin{cases}t^{m}P_{x},&\quad t\neq 0\\ P_{0,x},&\quad t=0\\ \end{cases}$
is smooth. Thus the Connes-Thom isomorphism $\mu_{t=0}$ transforms the $KK$-class $[P_{0,x}]$ to $[P_{x}]$. So the inverse map $\mu_{t=0}^{-1}$ transforms from $[P_{x}]$ to $[P_{0,x}]$. The isomorphism $\mu_{s=0}$ is induced by the splitting isomorphism from $T_{x}M$ to $gr(H)_{x}$ so it transforms to an operator $P^{\prime}_{x}$ on $T_{x}M$. As a result, we attach these operators over $x\in M$ and obtain a family of operators $P^{\prime}=P^{\prime}_{x}$ such that the product with $[\mathcal{E}]$ transforms $[\sigma_{H}(P)]\in KK_{G}(\mathbb{C},C^{\ast}(Gr(H)))$ to $[P^{\prime}]\in KK_{G}(\mathbb{C},C^{\ast}(TM))$. Take the Fourier transform from $C^{\ast}(TM)$ to $C_{0}(T^{\ast}M)$ we obtain $\sigma^{\prime}(P)(x,\xi)$.
Note that for a differential operator $P=\sum_{\lvert\alpha\rvert_{H}\leq m}c_{\alpha}X^{\alpha}$ on $M$, the term $P_{0,x}$ in above process is $P_{0,x}=\sum_{\lvert\alpha\rvert_{H}=m}c_{\alpha}[X]_{x}^{\alpha}\in\mathcal{U}(gr(H)_{x})$, and the $KK$-class only depend on the principal part, so we can change the splitting isomorphism to the map $[X]_{x}\in gr(H)_{x}\mapsto X_{x}\in T_{x}M$. Thus we can easily see that $P^{\prime}_{x}$ can be taken to be the freezing coefficient of the H-principal component of the original operator $P$.
Take the functional calculus by
$T(a)=\begin{pmatrix}0&a^{\ast}/(1+aa^{\ast})\\ a/(1+a^{\ast}a)&0\end{pmatrix}$
on both $P$ and $\sigma^{\prime}(P)$, we see that the same result holds replacing $P$ by its bounded transformation. ∎
Finally, we conclude the index classes and symbol classes in different cases above, and lead to a basic expectation for transversally H-elliptic operators. The construction of symbol and index classes in $KK$-homology, and their correlation to the analytical and topological index of the operator on a compact manifold are summarized as follows:
  • The index class, expressed as the $KK$-pairing $[(L^{2}(E),P)]\in KK_{G}(A,\mathbb{C})$, with different preference of the $C^{\ast}$-algebra $A$ in varying cases. The index class represents the analytical index for operators.
  • The symbol class is represented as the $KK$-pairing $[(C(M;E)\otimes_{C(M)}B,\sigma(P))]\in KK_{G}(\mathbb{C},B)$, where $B$ is selected to be the symbol algebra of “negative order” terms in different cases. It is constructed from the symbol depending on the topological properties of $P$ and $M$, such symbol class can be understood as the topological index. Especially, in the transversally elliptic case, the “negative order” algebra $S_{tr}(M)$ is replaced by the $KK$-equivalent Clifford algebra $Cl_{\Gamma}(TM)=C_{0}(TM)\otimes_{C(M)}Cl_{\Gamma}(M)$.
Thus, in the transversally H-elliptic case defined in the following section, we expect to construct some appropriate algebras $A$, $B$ and construct the symbol class and index class in a similar fashion:
$[P]\in KK_{G}(A,\mathbb{C}),\ [\sigma_{H}(P)]\in KK_{G}(\mathbb{C},B).$
Chapter 4 Construction for transversally H-elliptic operators
In this chapter, we explains the definition of transversally H-elliptic operators.
4.1 Basic notions
To articulate the concept of transversally H-elliptic operators, preliminary notions related to the transversally elliptic operators are required.
Let $M$ be a compact filtered Riemannian manifold equipped with a smooth, isometric action of a Lie group $G$. Recall that for each point $x\in M$, we have the mapping $f_{x}:G\to M$, $g\mapsto g(x)$. The derivative of $f_{x}$ at $e\in G$ is denoted by $f_{x}^{\prime}:\mathfrak{g}\to T_{x}M$.
Definition 4.1.1.
Define $\psi_{x}:T_{x}M\to gr(H)_{x}$ to be the orthogonal bundle map, which maps $v=v_{0}+v_{1}+\dots+v_{n}\in T_{x}M$, where all $v_{i},v_{j}$ are orthogonal in the Riemannian structure of $M$ and $v_{i}\in H^{i}$ for each $i$, to
$[v]:=[v_{0}]+[v_{1}]+\dots+[v_{n}]\in gr(H)_{x}=\bigoplus_{i=1}^{n}H^{i}_{x}/H^{i-1}_{x}.$
We set
$h_{x}:=\psi_{x}\circ f_{x}^{\prime}:\mathfrak{g}\to T_{x}M\to gr(H)_{x}.$
It induces a $G$-invariant quadratic form $\tilde{q}$ on $\xi\in gr(H)_{x}^{\ast}$ by
$\tilde{q}_{x}(\xi)=\lVert h_{x}^{\ast}(\xi)\rVert^{2},$
and we obtain a $G$-invariant map $\tilde{\varphi}_{x}=h_{x}h_{x}^{\ast}:gr(H)^{\ast}_{x}\to gr(H)^{\ast}_{x}$.
Remark 4.1.2.
Note that $\xi\in gr(H)^{\ast}_{x}$ corresponds to covectors that are orthogonal to the orbit passing through $x$ if and only if $\tilde{q}_{x}(\xi)=0$, or if and only if $\tilde{\varphi}_{x}(\xi)=0$.
Let $Y_{1},\dots,Y_{n}$ be an orthonormal basis for $\mathfrak{g}$ and the same notation is used for the corresponding left-invariant vector fields on $G$. Let $\tilde{Y_{1}},\dots,\tilde{Y_{n}}$ be the corresponding first-order differential operators (Lie derivatives by vector fields on $M$ corresponding to $Y_{i}$’s).
Definition 4.1.3.
The orbital Laplacian operator $\Delta_{G}$ with ordinary symbol $\sigma(\Delta_{G},x,\xi)={q}_{x}(\xi)$ is defined by
$\Delta_{G}:=\sum_{i}-\tilde{Y_{i}^{2}}.$
To define transversally H-elliptic operators from the perspective of the tangent Clifford symbol class, we first define the space of Clifford sections $Cl_{\tilde{\Gamma}}(M)$. Recall that we have defined $Cl_{V}(M)$ in Definition 2.2.4 for any vector bundle $V$.
Definition 4.1.4.
Define the vector bundle
$\tilde{\Gamma}:=\{h(s)\ |\ s\in C_{0}(M;M\times\mathfrak{g})\}.$
The map $f^{\prime}$ in Definition 2.2.4 is replaced by the $h$ map.
Then $Cl_{\tilde{\Gamma}}(M)$ is the algebra of continuous sections of the sub-algebra of $\text{Cliff}(gr(H))$ generated by the vectors tangent to the orbit direction.
4.2 Construction of leaf-wise Dirac operator
In this section, we give the construction of the orbital Dirac operator with respect to the $G$-action on $M$ in [kasparov2022k]. The construction can be derived by the following three steps:
  • For each $x\in M$, we first construct a canonical differential Dirac operator on the quotient $G/M_{x}$, where $M_{x}$ is the stability subgroup of $G$ at $x\in M$. The Lie algebra of the stability subgroup $M_{x}$ will be denoted $\mathfrak{m}_{x}$.
    Lemma 4.2.1 ([kasparov2022k]).
    The space of $L^{2}$-sections of the vector bundle $\Lambda^{\ast}(G/M_{x})$ is isomorphic to the $M_{x}$-invariant subspace $(L^{2}(G)\otimes\Lambda^{\ast}(\mathfrak{m}^{\bot}_{x}))^{M_{x}}$, where $M_{x}$ acts on $L^{2}(G)$ by right translations and $\Lambda^{\ast}$ means the exterior algebra.
    Definition 4.2.2.
    For any $v\in\mathfrak{g}$, let $g_{v}(t)$ be the one-parameter subgroup of $G$ corresponding to the vector $v$. The left-invariant differential operator $\partial/\partial v$ on smooth elements of $L^{2}(G)$ is defined to be the infinitesimal right translation by $g_{v}(t)$ on $L^{2}(G)$.
    Remark 4.2.3.
    Note that when $v\in\mathfrak{m}_{x}$, the differential operator $\partial/\partial v$ is 00 on $C_{c}^{\infty}(G/M_{x})$ and also on $(C_{c}^{\infty}(G)\otimes\Lambda^{\ast}(\mathfrak{m}^{\bot}_{x}))^{M_{x}}$.
    Choose any basis $\{\tilde{v}_{k}\}$ in $\mathfrak{m}^{\bot}_{x}$ and the dual basis $\{v_{k}\}$ in $\mathfrak{g}/\mathfrak{m}_{x}$. Define the differential Dirac operator on $(C_{c}^{\infty}(G)\otimes\Lambda^{\ast}(\mathfrak{m}^{\bot}_{x}))^{M_{x}}$ by the formula
    $\mathcal{D}_{G/M_{x}}=-i\sum_{k}(\text{ext}(\tilde{v}_{k})+\text{int}(\tilde{v}_{k}))\frac{\partial}{\partial v_{k}}.$
    Lemma 4.2.4 ([kasparov2022k]).
    For all $x\in M$, $\mathcal{D}_{G/M_{x}}$ are elliptic, and their principal symbols are given by $\sigma_{\mathcal{D}_{G/M_{x}}}=\text{ext}(\eta)+\text{int}(\eta)$.
  • Then we construct the Dirac operator acting on sections of the vector bundle $\Lambda^{\ast}(\mathcal{O}_{x})$ for each orbit $\mathcal{O}_{x}\cong G/M_{x}$. Using the homeomorphisms $f_{x}:G/{M_{x}}\cong\mathcal{O}_{x}$, we can push forward all these operators $\mathcal{D}_{G/M_{x}}$ to the orbits $\mathcal{O}_{x}$ in $M$. The operator $\mathcal{D}$ on the orbit $O$ will be denoted $\mathcal{D}_{O}$. By the construction above, it depends only on the orbit, not on the point $x$ of the orbit.
  • Using the embedding $\mathfrak{m}^{\bot}_{x}\cong T_{x}(O_{x})\subset T_{x}(M)$, one can redefine the operators $\mathcal{D}_{O_{x}}$ as operators acting on the bundles $\Lambda^{\ast}(M)|_{O_{x}}$. This will give the operators $\mathcal{D}_{O}$ on sections of the bundle $\Lambda^{\ast}(M)$ over the orbit $O$.
    Definition 4.2.5.
    The orbital Dirac operator $\mathcal{D}_{G}$ on $\Lambda^{\ast}(M)$ is formed by the family of operators $\{\mathcal{D}_{O_{x}}\}$ on $\Lambda^{\ast}(M)|_{O_{x}}$.
The above statements can be concluded as follows:
Proposition 4.2.6.
There exist an orbital Dirac operator $\mathcal{D}_{G}$ on $\Lambda^{\ast}(M)$ whose ordinary principle symbol is $\text{ext}(\varphi_{x}(\xi))+\text{int}(\varphi_{x}(\xi))$.
Proof.
The existence has been verified already. The ordinary principle symbol can be calculated by
$\sigma({\mathcal{D}_{G}})(x,\xi)$ $=\sum_{k}(\xi,f_{x}^{\prime}(v_{k}))(\text{ext}(f_{x}^{\prime}(v_{k}))+\text{int}(f_{x}^{\prime}(v_{k})))$
$=\sum_{k}({f_{x}^{\prime}}^{\ast}(\xi),v_{k})(\text{ext}(f_{x}^{\prime}(v_{k}))+\text{int}(f_{x}^{\prime}(v_{k})))$
$=\text{ext}(\varphi_{x}(\xi))+\text{int}(\varphi_{x}(\xi)).$
4.3 Adjoint in different viewpoints
The H-principle cosymbol $\sigma_{H}(\mathcal{D}_{G},x)$ at one point $x$ should be an operator on $\mathcal{G}=Gr(H)_{x}$. It defines an operator on $C_{c}^{\infty}(\mathcal{G})$ and extends to $C^{\ast}(\mathcal{G})$. The main purpose of this and the next section is to show the essential self-adjointness of
$\sigma_{H}(\mathcal{D}_{G},x)=-i\sum_{k}X_{k}(\text{ext}(e_{k})+\text{int}(e_{k})),$
where $X_{k}$ are elements in $gr(H)_{x}$ corresponding to generators of $\mathfrak{g}$ of highest order in the filtration structure and $e_{k}$ are the dual to $X_{k}$.
Note that we can define the adjoint operator of a left-invariant differential operator $P=\sigma_{H}(\mathcal{D}_{G},x)$, for example, in the sense of $C^{\ast}$-module morphism or in the sense of universal algebra. However, it turns out that these adjoint maps $P\to P^{\ast}$ are in fact the same.
Definition 4.3.1.
There are three definitions for the adjoint of an element in $P\in\mathcal{U}(gr(H)_{x})$:
  • (a)
    $P$ is identified as an element in the universal algebra $\mathcal{U}(gr(H)_{x})$. On the universal algebra $\mathcal{U}(gr(H)_{x})$, there is an involution map
    $-{id}:gr(H)_{x}\to gr(H)_{x}$
    which maps all $X_{i}$ to $-X_{i}$. It extended to get an anti-automorphism
    $i:\mathcal{U}(gr(H)_{x})\to\mathcal{U}(gr(H)_{x}).$
    The image $P^{\ast}:=i(P)$ is called the adjoint of $P$. This construction fits with all the algebraic structures we want in H-elliptic theory and satisfies
    $\sigma_{H}(D^{\ast})=\sigma_{H}(D)^{\ast}$
    for a differential operator $D$ on $M$.
  • (b)
    $P$ is identified as a morphism of the Hilbert $C^{\ast}(Gr(H)_{x})$-module $E=C^{\ast}(Gr(H)_{x})$ itself with inner product $\langle a,b\rangle=a^{\ast}b$. Then we can define the adjoint $P^{\ast}$ by
    $\langle Pf,g\rangle=\langle f,P^{\ast}g\rangle$
    by the dense range of $P$.
  • (c)
    $P$ is identified as a left-invariant differential operator, which can be transformed to an essentially homogeneous distribution on $Gr(H)_{x}$ by $(u,f)=(Pf)(e)$ for $f\in C_{c}^{\infty}(Gr(H)_{x})$. It corresponds to an operator $Op(u)$ by
    $(Op(u)(f))(x)=(u,f\circ l_{x}).$
    By the left-invariance of $P$, we see that $Op(u)=P$. The involution operator is defined by
    $(u^{\ast},f):=\overline{(u,f^{\ast})}.$
    Such involution on $u$ induces a adjoint $P^{\ast}$ for the corresponding operator $P=Op(u)$ of $u$.
Theorem 4.3.2.
The above three definitions for the adjoint operator are equivalent.
Proof.
We prove for a general nilpotent Lie group $\mathcal{G}$. We consider only unimodular groups, which means that the Haar measure is both left- and right-invariant.
Recall that we have defined the convolution and involution maps on $C^{\infty}_{c}(\mathcal{G})$ in Definition 1.3.2. Let $f,g\in C_{c}^{\infty}(\mathcal{G})$, the convolution $f\ast g$ is defined by
[…]
The involution $\ast$-operator is defined by
$f^{\ast}:=\overline{f(t^{-1})}.$
On $C^{\ast}(\mathcal{G})$-module $E=C^{\ast}(\mathcal{G})$, the product is given by
$\langle f,g\rangle=f^{\ast}\ast g.$
$(a)\Leftrightarrow(b)$: It suffices to prove that $\langle Xf,g\rangle=\langle f,-Xg\rangle$ for any first-order differential operator $X\in\mathfrak{g}$ and $f,g\in C_{c}^{\infty}(\mathcal{G})$. It can be proved by direct calculation that
[…]
$=\int_{G}X(\overline{f(y)}g(yx)))dy$
$=0$
by the invariance of the Haar measure.
$(a)\Leftrightarrow(c)$: Assume $f\in C_{c}^{\infty}(\mathcal{G})$. It suffice to prove that for $u$ corresponds to $X\in\mathfrak{g}$ by $(u,f)=Xf(e)$, $Op(u^{\ast})$ transform $f$ to $-Xf$. $Op(u^{\ast})f$ maps $x\in\mathcal{G}$ to
$\overline{X(\bar{f}\circ l_{x}\circ\iota)}(e)$
$=$ $dl_{x}\circ d\iota_{e}\circ\overline{X(\bar{f}(\cdot)})(e)$
$=$ $dl_{x}\circ(-X)(f(\cdot))(e)$
$=$ $-X(f)(x).$
The second equation holds because $d\iota_{e}=-\text{id}_{T_{e}\mathcal{G}}$.
Thus the adjoint constructions in $(a),(b),(c)$ are equivalent.
4.4 Essential self-adjointness of $\sigma_{H}(\mathcal{D}_{G})$
Let $\mathcal{D}_{G}$ be the orbital Dirac operator on $\Lambda^{\ast}(M)$. For each $x\in M$, take the $C^{\ast}(Gr(H)_{x})$-module $E=C^{\ast}(Gr(H)_{x})\otimes\Lambda^{\ast}_{x}(M)$. Then the H-principle cosymbol of $\mathcal{D}_{G}$ is an unbounded operator on the Hilbert module $E$
$\sigma_{H}(\mathcal{D}_{G},x):C^{\ast}(Gr(H)_{x})\otimes\Lambda^{\ast}_{x}(M)\to C^{\ast}(Gr(H)_{x})\otimes\Lambda^{\ast}_{x}(M).$
In this section, we want to prove the following theorem:
Theorem 4.4.1.
For each $\lambda\geq 1$ large enough, the image of $\sigma_{H}(\mathcal{D}_{G},x)\pm i\lambda$ is dense in the algebra $C^{\ast}(Gr(H)_{x})\otimes\Lambda^{\ast}_{x}(M)$.
Thus $\sigma_{H}(\mathcal{D}_{G},x)$ is a regular operator as a morphism on the $C^{\ast}(Gr(H)_{x})$-module $E=C^{\ast}(Gr(H)_{x})\otimes\Lambda^{\ast}_{x}(M)$.
By the formal adjoint argument in the last section,
$\sigma_{H}(\mathcal{D}_{G},x)=-i\sum_{k}X_{k}(\text{ext}(e_{k})+\text{int}(e_{k}))$
is formally self-adjoint, where $X_{k}$ are elements in $gr(H)_{x}$ corresponding to generators of $\mathfrak{g}$ of highest order in the filtration structure and $e_{k}$ are the dual to $X_{k}$.
Lemma 4.4.2.
Let $D$ be a first-order formally self-adjoint differential operator on $Gr(H)_{x}$ which has finite propagation speed. Namely, $c_{D}<\infty$, where
$c_{D}=\sup_{z\in{Gr(H)_{x}}}\sup_{\xi\in T_{z}^{\ast}{Gr(H)_{x}},\lVert\xi\rVert=1}\lVert\sigma_{D}(z,\xi)\rVert.$
Then, $D$ is essentially self-adjoint on the Hilbert space $L^{2}(Gr(H)_{x})\otimes_{\mathbb{C}}\Lambda_{x}^{\ast}(M)$.
Proof.
First, for each compact subset $K\subset Gr(H)_{x}$, there exists a family of Friedrich’s mollifiers $L^{2}(K)\otimes V\to L^{2}(Gr(H)_{x})\otimes V$, which is a sequence of bounded operators with the following properties:
  • (i)
    each $F_{n}$ is a contraction;
  • (ii)
    the image of each $F_{n}$ is contained in $C_{c}^{\infty}({Gr(H)_{x}})\otimes{\Lambda^{\ast}_{x}(M)}$;
  • (iii)
    for all $v\in L^{2}({Gr(H)_{x}})\otimes{\Lambda^{\ast}_{x}(M)}$, $F_{n}v\to v$ and $F_{n}^{\ast}v\to v$ in the $L^{2}({Gr(H)_{x}})$-norm as $n\to\infty$;
  • (iv)
    for any first-order differential operator $D$, sequence $[D,F_{n}]=[D,F_{n}^{\ast}]$ of operators on $C_{c}(K)\otimes V$ is uniformly bounded in operator norm.
[…]
$F_{n}=F^{0}_{n}\otimes 1$ $:L^{2}(K)\otimes{\Lambda^{\ast}_{x}(M)}\to L^{2}(\mathbb{R}^{d})\otimes{\Lambda^{\ast}_{x}(M)},$
[…]
Cover $K\subset Gr(H)_{x}$ by finitely many coordinate charts, and use a finite smooth partition of unity consisting of compactly supported functions to patch together mollifiers constructed as above.
To prove the essential self-adjointness of $D$, we need to show that for $v,w\in L^{2}({Gr(H)_{x}})\otimes{\Lambda^{\ast}_{x}(M)}$ such that for all $u\in C_{c}^{\infty}({Gr(H)_{x}})\otimes{\Lambda^{\ast}_{x}(M)}$,
$\langle Du,v\rangle=\langle u,w\rangle.$
(say $v$ is in the maximal domain of $D$ if there is a $w$ satisfying this condition), there exists a sequence $v_{n}$ in $C_{c}^{\infty}({Gr(H)_{x}})\otimes{\Lambda^{\ast}_{x}(M)}$ converging to $v$ in $L^{2}({Gr(H)_{x}})\otimes{\Lambda^{\ast}_{x}(M)}$ and such that $Dv_{n}$ converges to $w$ in $L^{2}({Gr(H)_{x}})\otimes{\Lambda^{\ast}_{x}(M)}$ (say $v$ is in the minimal domain of $D$ if there is a $w$ satisfying this condition).
Assume first that $v$ is compactly supported in some $K\subset Gr(H)_{x}$, let $(F_{n})$ be a family of Friedrich’s mollifiers for $K$. Consider the sequence $F_{n}v$, which is in $C_{c}^{\infty}({Gr(H)_{x}})\otimes{\Lambda^{\ast}_{x}(M)}$ and converges in $L^{2}({Gr(H)_{x}})$-norm to $v$. For any $u\in C_{c}^{\infty}({Gr(H)_{x}})\otimes{\Lambda^{\ast}_{x}(M)}$,
$\langle DF_{n}v,u\rangle$ $=\langle v,F_{n}^{\ast}Du\rangle$
$=\langle v,[F_{n}^{\ast},D]u\rangle+\langle v,DF_{n}^{\ast}u\rangle$
$=\langle v,[F_{n}^{\ast},D]u\rangle+\langle w,F_{n}^{\ast}u\rangle.$
Hence $\lVert DF_{n}v\rVert\leq\lVert[F_{n}^{\ast},D]\rVert\lVert v\rVert+\lVert w\rVert$ for all $n$, the sequence $(DF_{n}v)$ is uniformly bounded, thus has a convergent subsequence and still denote it by $(DF_{n}v)$. Since
$\langle u,DF_{n}v\rangle=\langle Du,F_{n}v\rangle\to\langle Du,v\rangle=\langle u,w\rangle,$
the limit is $w$.
By Mazur’s lemma, there is a sequence $(v_{n})$ in $C_{c}^{\infty}({Gr(H)_{x}})\otimes{\Lambda^{\ast}_{x}(M)}$ consisting of convex combinations of the sequence, such that $v_{n}\to v$ and $F_{n}v\to w$ which strongly converges to $w$ as $n\to\infty$. Thus we have proved that when $v$ in the maximal domain is compactly supported, $v$ is in the minimal domain of $D$.
It is obvious that ${Gr(H)_{x}}$ is complete and simply connected and $D$ has finite propagation speed $c_{D}=\sup_{z\in{Gr(H)_{x}}}\sup_{\xi\in T_{z}^{\ast}{Gr(H)_{x}},\lVert\xi\rVert=1}\lVert\sigma_{D}(z,\xi)\rVert$.
Let $(f_{n}:\mathbb{R}\to[0,1])$ be a sequence of smooth, compactly supported functions such that $f(t)=1$ for all $t\leq n$, and such that the sequence $(\sup_{t\in\mathbb{R}}|f_{n}^{\prime}(t)|)$ of real numbers tends to zero as $n$ tends to infinity. Fix $z_{0}\in{Gr(H)_{x}}$, and define
$g_{n}(z):=f_{n}(d(z,z_{0})).$
Then $g_{n}$ is smooth and compactly supported, the sequence $\sup_{z\in Gr(H)_{x}}\lVert dg_{n}(z)\rVert_{n=0}^{\infty}$ tends to zero as $n$ tends to infinity, and that $(g_{n})$ tends to one uniformly on compact subsets of ${Gr(H)_{x}}$.
For general elements $v,w\in L^{2}({Gr(H)_{x}})\otimes{\Lambda^{\ast}_{x}(M)}$ with the condition that $v$ is in the maximal domain, let $v_{n}=g_{n}v$. Then
$\langle{D}u,v_{n}\rangle$ $=\langle g_{n}{D}u,v\rangle$
$=\langle[g_{n},{D}]u,v\rangle+\langle{D}g_{n}u,v\rangle$
$=\langle\sigma_{D}(dg_{n})u,v\rangle+\langle g_{n}u,w\rangle$
$=\langle u,\sigma_{D}(dg_{n})v+g_{n}w\rangle$
for all $u\in C_{c}^{\infty}({Gr(H)_{x}})\otimes{\Lambda^{\ast}_{x}(M)}$. So $v_{n}$ is in the maximal domain of ${D}$ for all $n$. Since $v_{n}$ is compactly supported by compact support of $g_{n}$, $v_{n}$ is in the minimal domain of ${D}$. So we can apply the closure of ${D}$ on $v_{n}$ and for any $u\in C_{c}^{\infty}({Gr(H)_{x}})\otimes{\Lambda^{\ast}_{x}(M)}$,
$\langle{D}v_{n},u\rangle$ $=\langle v,g_{n}{D}u\rangle$
$=\langle v,[g_{n},{D}]u\rangle+\langle v,{D}g_{n}u\rangle$
$=\langle v,\sigma_{D}(dg_{n})u\rangle+\langle w,g_{n}u\rangle,$
and
$|\langle v,[g_{n},{D}]u\rangle|$ $\leq c_{D}\sup_{z\in Gr(H)_{x}}\lVert dg_{n}(z)\rVert_{n=0}^{\infty}\lVert v\rVert_{L^{2}}\lVert u\rVert_{L^{2}}\to 0$
as $n\to\infty$.
$\lim_{n\to\infty}\langle{D}v_{n},u\rangle=\lim_{n\infty}\langle w,g_{n}u\rangle=\langle w,u\rangle.$
Since $u$ is arbitrary, $({D}v_{n})$ converges to $w$. So $v$ is in the minimal domain of ${D}$. So the maximal domain of $D$ is contained in the minimal domain of $D$, thus $D$ is essentially self-adjoint. ∎
For any $\kappa,\lambda\in\mathbb{R},\lambda\neq 0$, let
$T_{\kappa,\lambda}:=\sigma_{H}(\mathcal{D}_{G},x)+i\kappa\sigma_{\sigma_{H}(\mathcal{D}_{G},x)}(z,d\rho(z))+i\lambda.$
Corollary 4.4.3.
Assume that $2|\kappa|\cdot\lVert c_{\sigma_{H}(\mathcal{D}_{G},x)}\rVert<|\lambda|$, then $T_{\kappa,\lambda}(C_{c}^{\infty}(Gr(H)_{x})\otimes\Lambda^{\ast}_{x}(M))$ is dense in $L^{2}(Gr(H)_{x})\otimes\Lambda_{x}^{\ast}(M)$.
Moreover, there is an inverse $T_{\kappa,\lambda}^{-1}$ which is a bounded operator on $L^{2}(Gr(H)_{x})\otimes\Lambda_{x}^{\ast}(M)$.
Proof.
Apply to Lemma 4.4.2 with $D=\sigma_{H}(\mathcal{D}_{G},x)+i\kappa\sigma_{\sigma_{H}(\mathcal{D}_{G},x)}(z,d\rho(z))$, it is formally self-adjoint and has finite propagation speed since $\sigma_{H}(\mathcal{D}_{G},x)$ has finite propagation speed. So $D$ is essentially self-adjoint and thus $D\pm\lambda i$ has dense image in $L^{2}(Gr(H)_{x})\otimes\Lambda_{x}^{\ast}(M)$ for $\lambda$ large enough. By direct calculation,
$\lVert T_{\kappa,\lambda}(u)\rVert^{2}\leq(\lambda^{2}-2|\kappa\lambda|\cdot\lVert c_{D}\rVert)\cdot\lVert u\rVert^{2}.$
So the operator $T_{\kappa,\lambda}^{-1}$ is a bounded on $L^{2}(Gr(H)_{x})\otimes\Lambda_{x}^{\ast}(M)$. ∎
Proof.
(of Theorem 4.4.1) Denote the closure of $\sigma_{H}(\mathcal{D}_{G},x)$ with $L^{2}$-inner product by $D_{1}$ and the closure with $C^{\ast}$-inner product by $D_{2}$.
Choose $\kappa$ positive and large enough so that the function of $z\in Gr(H)_{x}$, $e^{-\kappa\rho(z)}$ is in $L^{2}(Gr(H)_{x})$, where $\rho(z)$ the distance function from the identity $e$ to $z\in Gr(H)_{x}$.
Let $v\in C_{c}^{\infty}(Gr(H)_{x})\otimes\Lambda^{\ast}_{x}(M)$. By Corollary 6.2.4, the operator $T_{0,\pm\lambda}=D_{1}\pm i\lambda$ and $T_{\kappa,\pm\lambda}=D_{1}\pm i\lambda$ are invertible on $L^{2}(Gr(H)_{x})\otimes\Lambda_{x}^{\ast}(M)$. So we can solve the equation $(D_{1}\pm i\lambda)(u)=v$ and $T_{\kappa,\pm\lambda}(u)=e^{\kappa\rho(z)}v$ in $L^{2}(Gr(H)_{x})\otimes\Lambda_{x}^{\ast}(M)$. By direct calculation $(D_{1}\pm i\lambda)(u-e^{-\kappa\rho(z)}u_{1})=0$, so $u=e^{-\kappa\rho(z)}u_{1}$. Now we have $e^{-\kappa\rho(z)}$ in $L^{2}(Gr(H)_{x})$ and $u_{1}$ lie in $L^{2}(Gr(H)_{x})\otimes\Lambda_{x}^{\ast}(M)$. Thus the product $u=e^{-\kappa\rho(z)}u_{1}$ lies in $L^{1}(Gr(H)_{x})\otimes\Lambda_{x}^{\ast}(M)$ and its $L^{1}$-norm is restricted by
$\lVert u\rVert_{L^{1}\otimes\Lambda_{x}^{\ast}(M)}\leq\lVert e^{-\kappa\rho(z)}\rVert_{L^{2}}\lVert u_{1}\rVert_{L^{2}\otimes\Lambda_{x}^{\ast}(M)}.$
Thus $u\in C^{\ast}(Gr(H)_{x})\otimes\Lambda_{x}^{\ast}(M)$.
Similarly, we see that
$\lVert(1-g_{n})u\rVert_{C^{\ast}\otimes\Lambda_{x}^{\ast}(M)}\leq\lVert(1-g_{n})u\rVert_{L^{1}\otimes\Lambda_{x}^{\ast}(M)}\leq\lVert e^{-\kappa\rho(z)}\rVert_{L^{2}}\lVert(1-g_{n})u_{1}\rVert_{L^{2}\otimes\Lambda_{x}^{\ast}(M)},$
where $g_{n}$ is defined in the proof of Lemma 4.4.2.
So when $n\to\infty$, we see that the elements $g_{n}u\in C_{c}(G)$ converges to $u$ in the $C^{\ast}(Gr(H)_{x})$-module $C^{\ast}(Gr(H)_{x})\otimes\Lambda_{x}^{\ast}(M)$.
Let $u_{n}=g_{n}\ast g_{n}u$, since $g_{n}u$ is in $L^{1}$ and compactly supported, $g_{n}$ is smooth, we have $u_{n}\in C_{c}^{\infty}(G)\otimes\Lambda_{x}^{\ast}(M)$. Since $g_{n}\to 1$ in $L^{1}$ and thus in $C^{\ast}$, $1-g_{n}$ is uniformly bounded and decreasing, and $g_{n}u\to u$ also in $L^{1}$ and thus in $C^{\ast}$ when $n\to\infty$, we see that $u_{n}$ converges to $u$ in $C^{\ast}(Gr(H)_{x})\otimes\Lambda_{x}^{\ast}(M)$.
Since $D_{2}=D_{1}=\sigma_{H}(\mathcal{D}_{G},x)$ on $C_{c}^{\infty}(G)\otimes\Lambda_{x}^{\ast}(M)$, we have
$(D_{2}\pm i\lambda)(u_{n})$
$=$ $(D_{1}\pm i\lambda)(u_{n})$
$=$ $g_{n}\ast D_{1}(g_{n}u)\pm i\lambda u_{n}$
$=$ $g_{n}\ast\sigma_{D_{1}}(x,dg_{n})u+g_{n}\ast g_{n}D_{1}u\pm g_{n}\ast g_{n}(i\lambda u)$
$=$ $g_{n}\ast\sigma_{D_{1}}(x,dg_{n})u+g_{n}\ast g_{n}v.$
Note that $\sigma_{D_{1}}(x,dg_{n})\to 0$,
$\lim_{n\to\infty}(D_{2}\pm i\lambda)(u_{n})=v.$
So the images of $D_{2}\pm i\lambda$ are dense in $C_{c}^{\infty}(Gr(H)_{x})\otimes\Lambda^{\ast}_{x}(M)$. As a result the images of $\sigma_{H}(\mathcal{D}_{G},x)\pm i\lambda$ are dense in the $C^{\ast}(Gr(H)_{x})\otimes\Lambda^{\ast}_{x}(M)$, i.e., $\sigma_{H}(\mathcal{D}_{G},x)$ is a regular operator as a morphism on the $C^{\ast}(Gr(H)_{x})$-module $E=C^{\ast}(Gr(H)_{x})\otimes\Lambda^{\ast}_{x}(M)$. ∎
4.5 Definition for transversally H-elliptic operators
From the basic idea of transversally elliptic theory, we shall define the transversal H-ellipticity by requiring operators to be H-elliptic in the directions transversal to $G$-orbits.
The fundamental definition is motivated by the tangent-Clifford symbol for transversally elliptic operators defined in Definition 2.2.5. Recall that for a transversally elliptic operator $P$, its tangent-Clifford symbol $[\sigma^{tcl}(P)]$ is in the $KK$-group $KK_{G}(\mathbb{C},Cl_{\Gamma}(TM))=KK_{G}(\mathbb{C},Cl_{\Gamma}(M)\otimes_{C(M)}C_{0}(TM))$.
Replacing the $C_{0}(TM)\cong C^{\ast}(TM)$ by $C^{\ast}(Gr(H))$ and $Cl_{{\Gamma}}(M)$ by $Cl_{\tilde{\Gamma}}(M)$ in the group $KK_{G}(\mathbb{C},Cl_{\Gamma}(TM))=KK_{G}(\mathbb{C},Cl_{\Gamma}(M)\otimes_{C(M)}C_{0}(TM))$, we intend to derive the formal formula of the tangent-Clifford symbol for operators that can induce a $KK$-class $[\sigma_{H}^{tcl}(P)]$ in
$KK_{G}(\mathbb{C},Cl_{\tilde{\Gamma}}(M)\otimes_{C(M)}C^{\ast}(Gr(H))).$
The definition of $KK$-classes requires $1-\sigma^{tcl}_{H}(P)^{2}$ to be a compact multiplier, which motivates our definition.
Definition 4.5.1.
Let $M$ be a compact filtered Riemannian manifold and $G$ be a Lie group which acts on $M$ isometrically and preserves the filtered structure. We call a pseudo-differential operator of 00-order $P$ is transversally H-elliptic if at each point $x\in M$, the multiplier term
$(1-\sigma_{H}^{2}(P,x))(1+\sigma_{H}(\Delta_{G},x))^{-1}\in M(E_{x}\otimes_{\mathbb{C}}C^{\ast}(Gr(H)_{x}))$
is compact as operators on $C^{\ast}(Gr(H)_{x})$-modules.
Remark 4.5.2.
Note that
  • In the multiplier term $(1-\sigma_{H}^{2}(P,x))(1+\sigma_{H}(\Delta_{G},x))^{-1}$, the first term $(1-\sigma_{H}^{2}(P,x))$ is a multiplier of $E_{x}\otimes C^{\ast}(Gr(H)_{x})$ by definition of H-cosymbol, and the second term $(1+\sigma_{H}(\Delta_{G},x))^{-1}$, by the essentially self-adjointness of $\sigma_{H}(\mathcal{D}_{G},x)$ as a $C^{\ast}(Gr(H)_{x})$ morphism in Theorem 4.4.1, also represent a multiplier of $C^{\ast}(Gr(H)_{x})$. When vector bundle $E$ is trivial (in general cases we take the local trivialization such that $\text{End}(E)$ part becomes a matrix), the condition in Definition 4.5.1 requires that for all $x\in M$, the product multiplier element $(1-\sigma_{H}^{2}(P,x))(1+\sigma_{H}(\Delta_{G},x))^{-1}$ lies in the $C^{\ast}$-algebra $C^{\ast}(Gr(H)_{x})$.
  • The leaf-wise sub-Laplacian operator $\Delta_{G}$ is employed instead of a leaf-wise H-elliptic operator, because the leaf-wise direction contains several different order terms if an additional estimate is not obtained. Thus, an appropriate operator like $\Delta_{G}$ with H-elliptic property cannot be readily presumed in the leaf-wise direction in general cases.
Remark 4.5.3.
If the group $G$ acts trivially on $M$, then the condition will reduce to $1-\sigma_{H}^{2}(P,x)$ is compact. In this case, $1-\sigma_{H}^{2}(P,x)$ is homogeneous up to smooth so in the pseudo-differential calculus, the compactness of a multiplier is equivalent to of negative order. The condition is equivalent to the existence of parametrix for $\sigma(P,x)$, which implies the Rockland condition at each point $x\in M$ and leads to the H-ellipticity, Conversely, it can be proved that the cosymbols of H-elliptic operators are unbounded Fredholm multipliers (refer to [joachim2003unbounded] for the definition). Considering the bounded transform expressed in Section 3.2, then $1-\sigma^{2}_{H}(P)$ is a compact multiplier.
The symbol class and index class can be defined for such transversally H-elliptic operators.
Definition 4.5.4.
Define the operator $\sigma_{H}^{tcl}(P)$ by
$\sigma^{tcl}_{H}(P):=\sigma_{H}(P)\hat{\otimes}1+(1-\sigma_{H}(P)^{2})^{1/2}\hat{\otimes}f_{\tilde{\Gamma}}\in\mathcal{L}(C(M;E)\hat{\otimes}_{C(M)}C^{\ast}(Gr(H))\otimes_{C(M)}Cl_{\tilde{\Gamma}}(M)).$
Here $f_{\tilde{\Gamma}}$ is a multiplier of $Cl_{\tilde{\Gamma}}(M)\otimes_{C(M)}C^{\ast}(Gr(H))$ defined as $\sigma_{H}(\mathcal{D}_{G})\otimes(1+\sigma_{H}(\Delta_{G}))^{-1/2}$. At each point, $\sigma_{H}(\mathcal{D}_{G},x)$ denotes an element in Clifford product space $\tilde{\Gamma}_{x}$ such that it represents a multiplier on $Cl_{\tilde{\Gamma}}(M)$, $(1+\sigma_{H}(\Delta_{G,x}))^{-1/2}$ is a multiplier on $C^{\ast}(Gr(H)_{x})$. Consider the tensor product over $C(M)$ to obtain the desired operator $f_{\tilde{\Gamma}}$.
By the construction of $\sigma_{H}^{tcl}(P)$, we have the following
Proposition 4.5.5.
$1-(\sigma^{tcl}_{H}(P))^{2}=(1-\sigma^{2}_{H}(P))(1-f_{\tilde{\Gamma}}^{2})=(1-\sigma_{H}^{2}(P))(1+\sigma_{H}(\Delta_{G}))^{-1}.$
Proof.
By direct calculation,
$1-(\sigma^{tcl}_{H}(P))^{2}=1-\sigma^{2}_{H}(P)\hat{\otimes}1-(1-\sigma_{H}(P)^{2})\hat{\otimes}f^{2}_{\tilde{\Gamma}}-(1-\sigma_{H}(P)^{2})^{1/2}[\sigma_{H}(P)\hat{\otimes}1,1\hat{\otimes}f_{\tilde{\Gamma}}].$
Since $\sigma_{H}(P)\hat{\otimes}1$ represents on $C^{\ast}(Gr(H))$ part and $1\hat{\otimes}\sigma_{H}(\mathcal{D}_{G})$ represents on $Cl_{\tilde{\Gamma}}(M)$ part, the commutator $[\sigma_{H}(P)\hat{{\otimes}}1,1\hat{\otimes}\sigma_{H}(\mathcal{D}_{G})]=0$. Let $f(a)={a}/{(1+a^{2})}$, then $1\hat{\otimes}f_{\tilde{\Gamma}}$ is the functional calculus by $f$ on $1\hat{\otimes}\sigma_{H}(\mathcal{D}_{G})$. As a result, $[\sigma_{H}(P)\hat{\otimes}1,1\hat{\otimes}f_{\tilde{\Gamma}}]=0$ and
$1-(\sigma^{tcl}_{H}(P))^{2}=$ $1-\sigma^{2}_{H}(P)\hat{\otimes}1-(1-\sigma_{H}(P)^{2})\hat{\otimes}f^{2}_{\tilde{\Gamma}}$
$=$ $(1-\sigma^{2}_{H}(P))(1-f_{\tilde{\Gamma}}^{2})$
$=$ $(1-\sigma_{H}^{2}(P))(1+\sigma_{H}(\Delta_{G}))^{-1}.$
Thus $1-(\sigma_{H}^{tcl}(P))^{2}$ is compact by definition of transversal H-ellipticity. So we can define that:
Definition 4.5.6.
Let $P$ be a transversally H-elliptic operator. The symbol class $[\sigma_{H}^{tcl}(P)]\in K^{G}_{0}(Cl_{\tilde{\Gamma}}(M)\otimes_{C(M)}C^{\ast}(Gr(H)))$ is defined by the $KK$-pairing
$[\sigma_{H}^{tcl}(P)]:=(C(M;E)\hat{\otimes}_{C(M)}C^{\ast}(Gr(H)){\otimes}Cl_{\tilde{\Gamma}}(M),\sigma^{tcl}_{H}(P)).$
The index class is given as follows:
Theorem 4.5.7.
Let $P$ be a transversally H-elliptic operator. Then $(L^{2}(E),P)$ defines an element in the group $K^{0}(C^{\ast}(G))$. It is denoted $[P]$ and called the index class of $P$.
Proof.
Let $\{v_{k}\}$ be a basis for the Lie algebra $\mathfrak{g}$ and $v_{k}^{\ast}$ the dual basis for $\mathfrak{g}$. Define the operator
$d_{G}:C_{c}^{\infty}(E)$ $\to C_{c}^{\infty}(E\otimes\mathfrak{g}_{M}^{\ast})$
$f$ $\mapsto\sum_{k}\frac{\partial f}{\partial v_{k}}\otimes v_{k}^{\ast},$
where $\frac{\partial}{\partial v_{k}}$ indicates the derivative along the one-parameter subgroup of $G$ corresponding to the vector $v_{k}$.
Let $K$ be the operator with the cosymbol $(1-\sigma_{H}^{2}(P))(1+\sigma_{H}(\Delta_{G}))^{-1}$ that is compact at each point $x\in M$.
Using Theorem 1.6.13, we see that the operator $K$ itself is compact. To prove that $(1-P^{2})e$ is compact for all $e\in C_{c}^{\infty}(G)$, let $F=d_{G}^{\ast}Kd_{G}$. As $\sigma_{H}(d_{G}^{\ast}d_{G})=\sigma_{H}(\Delta_{G})$, we prove that $\sigma_{H}(1-P^{2})-\sigma_{H}(F)$ is of negative order, such that $1-P^{2}$ is equal to $F$ modulo operators of negative order.
By Lemma 6.6 in [kasparov2017elliptic], for any arbitrary $e\in C_{c}(G)$, the representation of $e$ on $F$, $Fe=d_{G}^{\ast}Kd_{G}\circ e$ is equal to the product of the compact operator $d_{G}^{\ast}K$ and the bounded convolution operator with $d(e)$. Therefore, $Fe$ and thus $(1-P^{2})e$ is compact for arbitrary $e\in C_{c}(G)$, and by natural extension, for arbitrary $e\in C^{\ast}(G)$, establishing it as a $KK$-class in $K^{0}(C^{\ast}(G))$. ∎
Remark 4.5.8.
In general case that $M$ is locally compact, we ought to consider the index class in $K^{0}(C^{\ast}(G,C_{0}(M)))$ and symbol class in $\mathcal{R}KK(C_{0}(M),Cl_{\tilde{\Gamma}}(M)\otimes_{C(M)}C^{\ast}(Gr(H)))$ which is defined in [kasparov1988equivariant]. In the case that $M$ is compact, it can be simplified to our result by the induced map of the inclusion $\mathbb{C}\subset C(M)$ and the index class in $K^{0}(C^{\ast}(G))$ is called the distribution index of $P$.
Chapter 5 Index theorem for transversally H-elliptic operators
5.1 From elliptic to H-elliptic cases
In this section, we introduce the idea to prove the H-elliptic index theorem (Theorem 3.2.6), from the index theorem for elliptic operators.
Recall that the index theorem for elliptic operators (Theorem 2.1.4) is proved by using a rotation homotopy, which transforms multiplication variable $\xi$ in symbol $[\sigma(P)(x,\xi)]$ to the differential variable $\chi$ in the operator $P$. The Dolbeault element will be transformed into a $1$ element in $KK$-class. Such a homotopy induces a $KK$-product relation between $[P]$ and $[\sigma(P)]$ and leads to the index theorem.
Motivation for the proof
For an H-elliptic operator $P$ with conditions in theorem, in order to find such a relation represented by $KK$-product between the index class $[P]$ and symbol class $[\sigma_{H}(P)]$, it is most possible to find some similar homotopy to relate the representation elements of these two terms. However, there is a gap that $P$ is an operator on $M$ and $\sigma_{H}(P)$ is an operator on $Gr(H)$. The calculation for the H-cosymbol cannot be conducted on $TM$ or $T^{\ast}M$, preventing a direct computation of the H-cosymbol for both the symbol operator and the original operator in the same space, and thereby complicating the direct construction of a homotopy of rotation from the $\xi$-variable to $\chi$-variable. One possible way to do is to lift the operator $P$ to $Gr(H)$ and construct some homotopy on $Gr(H)$ to relate $P$ and $\sigma_{H}(P)$. Another possible way to analyze directly we found, is trying to analyze operators on the rectangle diagram
${Gr(H)}$${gr(H)}$${M}$${TM,}$
with the operator on $TM$ is the lift of $P$ to $TM$ and the operator on $Gr(H)$ is $\sigma_{H}(P)$, and embed them into this diagram. In this way, we could find some relation between the ordinary symbol $\sigma(P)$ and $\sigma(\sigma_{H}(P))$. However, $Gr(H)$ has no general coordinates, so we don’t know how to construct a natural homotopy on $Gr(H)$ or in such a diagram, and it’s difficult to deal with the Connes-Thom class directly because it can’t be represented explicitly. So we changed these approaches to the following, using the idea in the proof of the H-elliptic index theorem by Mohsen in [mohsen2022index]:
Main idea
We transform the symbol class from $K_{0}(C^{\ast}(Gr(H)))$ to $K^{0}(T^{\ast}M)$ by the composition of isomorphisms, see the composition $\mu_{s=0}\circ\mu_{t=0}^{-1}$ in the following diagram:
${K^{0}(gr(H)^{\ast})}$${K^{0}(T^{\ast}M)}$${K_{0}(C^{\ast}(Gr(H)))}$${\mathbb{Z}.}$$\scriptstyle{\mu_{t=0}}$$\scriptstyle{\mu_{s=0}}$$\scriptstyle{\mu_{t=1}}$$\scriptstyle{\mu_{s=1}}$
Note that it represent the $KK$-class $[\mathcal{E}]$, and $\mu_{t=0}$ is the Connes-Thom isomorphism. By Lemma 3.2.7, The $KK$-product of the symbol class $[\sigma_{H}(P)]$ and $[\mathcal{E}]$ can be represented by the Fourier transform of a family of operators $P^{\prime}$ on $T_{x}M$ with $x\in M$. In this way, we delete the Connes-Thom term and obtain a multiplication operator $\sigma^{\prime}(P)$ with coordinates $(x,\xi)$ on $TM$. Then we can do a similar way in the proof of Theorem 9.1 in [kasparov2022k] to show the existence of homotopy relating $\sigma^{\prime}(P)$ and the lifting of $P$ to $TM$.
We will use this approach to give a proof of Theorem 3.2.6 in detail. It will subsequently facilitate the derivation of the main theorem through a similar process.
Proof of Theorem 3.2.6
Proof.
Assume that $P_{0}$ is an H-elliptic differential operator and $P$ is its bounded transformation. The proof is separated into the following five parts:
  • (i)
    First, we come back to $C(M)$ in place of $\mathbb{C}$. We have the symbol class $\sigma_{H}(P)\in KK_{0}^{G}(C(M),C^{\ast}(Gr(H)))$ and the index class $[P]\in K^{0}(C^{\ast}(G,C(M)))$.
    Utilizing the Dolbeault element $[\mathcal{D}_{M}]\in K^{0}_{G}(C_{0}(TM))$ and the Bott–Dirac element $[\mathcal{C}]\in KK_{G}(C(M),C(M))$, we want to prove with $KK$-theoretic argument that $[\mathcal{C}]\otimes_{C(M)}[P]$ and $[\sigma_{H}(P)]$ $\otimes_{C^{\ast}(Gr(H))}[\mathcal{E}]\otimes_{C_{0}(TM)}[\mathcal{D}_{M}]$ yield identical elements, where $[\mathcal{E}]\in KK_{G}(C^{\ast}(Gr(H)),$ $C_{0}(TM))$ represents the Connes-Thom class we defined in section 3.2.
    The orthonormal coordinates of $T(TM)$ are regarded to be $(x,\xi,\chi,\zeta)$ with $(x,\xi)$, the coordinate of the manifold $TM$, $\chi$ the differential coordinate of $x$, and $\zeta$ the differential coordinate of $\xi$.
  • (ii)
    By the basic construction of $KK$-products, the $KK$-product element $[\mathcal{C}]\otimes_{C(M)}[P]$ can be represented by $(\mathcal{H},Q_{1}=N_{1}(\mathcal{C}\hat{\otimes}1+N_{2}1\hat{\otimes}\widetilde{P}))$, where =$\mathcal{H}=\{L^{2}(\Lambda^{\ast}(T_{x}M))\}_{x\in M}\otimes_{C(M)}L^{2}(E)$, $\widetilde{P}$ denotes a $KK$-theoretic connection of $P$. Note that $\mathcal{C}$ represents the Bott-Dirac operator on $L^{2}(\Lambda^{\ast}(T_{x}M))$, the operator $\widetilde{P}$ can be taken as the lift of $P$ to $TM$ and $1$ represents the identity operator on $L^{2}(\Lambda^{\ast}(M))$. The operator $Q_{1}$ denotes a pseudo-differential operator on $TM$.
    By Lemma 3.2.7, the $KK$-product $[\sigma_{H}(P)]\otimes_{C^{\ast}(Gr(H))}[\mathcal{E}]$ is represented by freezing coefficients of $\sigma_{H}(P)$ to the unit element $e\in Gr(H)_{x}$ for each $x\in M$ and transforming the family of the operators we acquired on $gr(H)_{x}$ to a family of operators on $T_{x}M$, and then taking the Fourier transform on such family. This process leads to the operator $\sigma^{\prime}(P)(x,\xi)$ on $C_{0}(TM)$, which is the Fourier transform of the freezing coefficient of the H-principal part of $P$ to each point $x\in M$. Consequently, the $KK$-product element $[\sigma_{H}(P)]\otimes_{C^{\ast}(Gr(H))}[\mathcal{E}]\otimes_{C_{0}(TM)}[\mathcal{D}_{M}]$ is represented as $(\mathcal{H},Q_{2}=N_{1}^{\prime}(\widetilde{\mathcal{D}_{M}}+N_{2}^{\prime}1\hat{\otimes}\sigma^{\prime}(P)))$. $1\hat{\otimes}\sigma^{\prime}(P)$ is considered over $TM$ with coefficient $(x,\xi)$, $\widetilde{\mathcal{D}_{M}}$ denotes the Dolbeault operator on $\{L^{2}(\Lambda^{\ast}(T_{x}M))\}_{x\in M}\otimes_{C(M)}L^{2}(E)$ lifted to $\mathcal{H}$, and the second $KK$-product is taken over $C_{0}(TM)$. Thus, the product operator $Q_{2}$ is a pseudo-differential operator on $TM$.
  • (iii)
    Now we construct the rotation homotopy and show that these two product operators are homotopic through this homotopy. It suffices to prove the existence of a uniform homotopy between operators $\mathcal{C}$ and $\mathcal{D}_{M}$, $\widetilde{P}$ and $\sigma^{\prime}(P)$.
    Note that $\widetilde{\mathcal{D}_{M}}$ and $\mathcal{C}\hat{\otimes}1$ are operators on $TM$, the symbol of Dolbeault operator is
    $\sigma_{\mathcal{D}_{M}}(x,\xi,\chi,\zeta)=\text{ext}(-\chi+i\zeta)+\text{int}(-\chi+i\zeta)$
    and the symbol of $\mathcal{C}$ is
    $\sigma_{\mathcal{C}}(x,\xi,\chi,\zeta)=\text{ext}(\xi+i\zeta)+\text{int}(\xi+i\zeta).$
    So the rotation homotopy in the $(\xi,\chi)$ variable:
    $\xi\mapsto\cos t\ \xi-\sin t\ \chi;\ \chi\mapsto\sin t\ \xi+\cos t\ \chi,\ t\in[0,\pi/2].$
    transforms $\mathcal{C}\hat{\otimes}1$ to $\widetilde{\mathcal{D}_{M}}$.
  • (iv)
    By the argument in Remark 1.4.6, the $\sigma^{\prime}(P)$ is given by the Fourier transform of the freezing coefficient operators of the H-principal part of $P$ to each $x\in M$. The rotation homotopy above, which transforms the multiplication with $\xi$-variable by differential with $x$ variable in the operator sense, transforms the multiplication operator $\sigma^{\prime}(P_{0})(x,\xi)$, which is a polynomial in $\xi$, to an operator with the same H-principal part with the differential operator $\tilde{P_{0}}$, and thus, by functional calculus, transforms $\sigma^{\prime}(P)(x,\xi)$ to an operator having the same $KK$-representation with $\tilde{P}$ on $TM$. As a result, up to lower order terms, $\sigma^{\prime}(P)$ is homotopic to $\tilde{P}$ by the rotation homotopy.
    During the homotopy we can choose appropriate $N_{1}^{t},N_{2}^{t}$ in the $KK$-product construction, such that it fits with the rotation homotopy and satisfies $N_{1}^{0}=N_{1},N_{1}^{\pi/2}=N_{1}^{\prime},N_{2}^{0}=N_{2},N_{2}^{\pi/2}=N_{2}^{\prime}$.
    Based on the above argument, we observed that there is a rotation homotopy between the operators representing $[\mathcal{C}]\otimes_{C(M)}[P]$ and $[\sigma_{H}(P)]\otimes_{C^{\ast}(Gr(H))}[\mathcal{E}]\otimes_{C_{0}(TM)}[\mathcal{D}_{M}]$ in $K^{0}_{G}(C(M))$. As a result, they represent the same $KK$-class.
  • (v)
    Following Remark 10.4 in [kasparov2022k], the restriction map from $KK_{G}^{0}(C(M))$ to $KK_{G}^{0}(\mathbb{C})$, induced by the inclusion $\mathbb{C}\subset C(M)$, assigns the index class $[P]$ to the analytical index of $P$. As $[\mathcal{C}]$ represents the trivial element $1\in KK_{G}(C(M),C(M))$, the result follows if we take the restriction of these two $KK$-classes to $KK^{0}_{G}(\mathbb{C})$ induced by the inclusion map.
That proves Theorem 3.2.6.
5.2 Main index theorem for transversally H-elliptic differential operators
Recall Definition 2.2.7 of the Clifford–Dolbeault element $[\mathcal{D}^{cl}_{M,{\Gamma}}]\in K^{0}(C^{\ast}(G,$ $Cl_{{\Gamma}}(TM)))$. The $KK$-element $[\mathcal{D}^{cl}_{M,\tilde{\Gamma}}]$ is defined similar to $[\mathcal{D}^{cl}_{M,{\Gamma}}]$.
Definition 5.2.1.
The Clifford Dolbeault element $[D^{cl}_{M,\tilde{\Gamma}}]\in K^{0}(C^{\ast}(G,Cl_{\tilde{\Gamma}}(TM)))$ is defined to be the $KK$-element $[(\mathcal{H},\pi,F)]$, where
  • $\mathcal{H}=L^{2}(\Lambda_{\mathbb{C}}^{\ast}(TM))$;
  • $F=\mathcal{D}_{M}$ is the same operator as the Dolbeault element $[\mathcal{D}_{M}]\in K^{0}_{G}(C_{0}(TM))$;
  • The representation $\pi$ of $C^{\ast}(G,Cl_{\tilde{\Gamma}}(TM))$ is induced from the action $\zeta\mapsto\text{ext}(\psi^{-1}(\zeta))+\text{int}(\psi^{-1}(\zeta))$ for $\zeta$ a section of $\tilde{\Gamma}$, where $\psi$ denotes the bundle isomorphism defined in Definition 4.1.1.
For a transversally H-elliptic operator $P$, we obtain the symbol class
$[\sigma^{tcl}_{H}(P)]\in K_{0}^{G}(C^{\ast}(Gr(H))\otimes_{C(M)}Cl_{\tilde{\Gamma}}(M)),$
and the index class
$[P]\in K^{0}(C^{\ast}(G)).$
The index theorem for transversally H-elliptic operators is presented as follows:
Theorem 5.2.2.
Let $M$ be a closed filtered Riemannian manifold and $G$ be a Lie group isometrically acting on $M$ and preserving the filtered structure. Let $P_{0}$ be a $G$-invariant differential operator on $M$ with its bounded transform
$P=\begin{pmatrix}0&P_{0}^{\ast}/(1+P_{0}P_{0}^{\ast})\\ P_{0}/(1+P_{0}^{\ast}P_{0})&0\end{pmatrix}$
being transversally H-elliptic of order 00, then we have
$[P]=j^{G}([\sigma^{tcl}_{H}(P)])\otimes_{C^{\ast}(G,C^{\ast}(Gr(H)))}j^{G}([\mathcal{E}])\otimes_{C^{\ast}(G,C_{0}(T^{\ast}M))}[\mathcal{D}_{M,\tilde{\Gamma}}^{cl}],$
where the $KK$-product on the right-hand side is considered
$KK(C^{\ast}(G),C^{\ast}(G,C^{\ast}(Gr(H))\otimes_{C(M)}Cl_{\tilde{\Gamma}}(M))$
$\otimes_{C^{\ast}(G,C^{\ast}(Gr(H))))}$ $KK(C^{\ast}(G,C^{\ast}(Gr(H))),C^{\ast}(G,C_{0}(T^{\ast}M)))$
$\otimes_{C^{\ast}(G,C_{0}(T^{\ast}M))}$ $KK(C^{\ast}(G,Cl_{\tilde{\Gamma}}(M)\otimes C_{0}(T^{\ast}M)),\mathbb{C}).$
The proof of the H-elliptic index theorem, Theorem 3.2.6 in Section 3.2, employed the approach in [kasparov2022k] by $KK$-theory. Subsequently, the proof for transversally H-elliptic cases parallels the methodology used in the proof of Theorem 10.2 in [kasparov2022k] for transversally elliptic operators by Kasparov.
5.3 Proof of the main theorem
First, we conclude the approach we used in the proof in section 5.1 as follows. We deformed the cosymbol $\sigma_{H}(P)$ to the operator $\sigma^{\prime}(P)(x,\xi)$ on $C_{0}(TM)$ by the Connes-Thom’s isomorphism. This allows both operators to reside on $TM$. Subsequently, a homotopy on symbol through rotation of coordinates on $T(TM)$ demonstrated that these two operators are homotopic.
Employing a similar approach for Theorem 3.2.6 allows us to give a proof for the main theorem (Theorem 5.2.2). We can try to transform the tangent-Clifford symbol into a symbol-like operator on $C_{0}(TM)$. Mainly, we need to prove that the two operators representing $KK$-classes $j^{G}([\sigma^{tcl}_{H}(P)]\otimes_{C^{\ast}(Gr(H)))}[\mathcal{E}])\otimes_{C^{\ast}(G,C_{0}(T^{\ast}M))}[\mathcal{D}_{M,\tilde{\Gamma}}^{cl}]$ and $[\mathcal{C}]\otimes_{C(M)}[P]$ are homotopic up to 00-elements in $KK$-classes.
Proof.
(of Theorem 5.2.2) Assume that $P_{0}$ is a transversally H-elliptic differential operator and $P$ is its bounded transformation. The proof is separated into the following five parts:
  • (i)
    Revert from $\mathbb{C}$ to $C(M)$ as in the proof of Theorem 3.2.6, the symbol class $[\sigma_{H}^{tcl}(P)]$ lies in $KK^{G}(C(M),C^{\ast}(Gr(H))\otimes_{C(M)}Cl_{\tilde{\Gamma}}(M)),$ and the index class $[P]$ lies in $KK(C^{\ast}(G,C(M)))$.
  • (ii)
    Take $\sigma^{\prime}(P)(x,\xi)$ to be the operator on $C_{0}(TM)$ we defined in Lemma 3.2.7. Then as in the proof of Theorem 3.2.6, we want to replace the class $[\sigma^{tcl}_{H}(P)]\otimes_{C^{\ast}(Gr(H)))}[\mathcal{E}]$ by some $[\sigma^{\prime}_{tcl}(P)]\in KK_{G}(C(M),C_{0}(T^{\ast}M)$. By definition, $\sigma^{tcl}_{H}(P)=\sigma_{H}(P)\hat{\otimes}1+(1-\sigma_{H}(P)^{2})^{1/2}\hat{\otimes}f_{\tilde{\Gamma}}$. The product over $C^{\ast}(Gr(H))$ with $[\mathcal{E}]$ transforms $\sigma_{H}(P)\hat{\otimes}1$ to $\sigma^{\prime}(P)\hat{\otimes}1$. Let $f_{\Gamma}$ is the ordinary cosymbol of the 00-order operator $\mathfrak{F}_{G}=\mathcal{D}_{G}(1+\mathcal{D}_{G}^{2})^{-1/2}$. Note that all the terms in $\mathcal{D}_{G}$ are normalized to 00-order, so the product with $[\mathcal{E}]$ transforms $f_{\tilde{\Gamma}}$ to $f_{{\Gamma}}$, and thus $(1-\sigma_{H}(P)^{2})^{1/2}\hat{\otimes}f_{\tilde{\Gamma}}$ to $(1-{\sigma^{\prime}(P)}^{2})^{1/2}\hat{\otimes}f_{{\Gamma}}$. The sum of the two resulting terms can be denoted by $\sigma^{\prime}_{tcl}(P)=\sigma^{\prime}(P)\hat{\otimes}1+(1-{\sigma^{\prime}(P)}^{2})^{1/2}\hat{\otimes}f_{{\Gamma}}$.
    Then we can represent the $KK$-product element $j^{G}([\sigma^{tcl}_{H}(P)])\otimes_{C^{\ast}(G,C^{\ast}(Gr(H)))}j^{G}([\mathcal{E}])$
    $\otimes_{C^{\ast}(G,C_{0}(T^{\ast}M))}[\mathcal{D}_{M,\tilde{\Gamma}}^{cl}]$ by the pairing
    $(L^{2}(E_{1}\hat{\otimes}_{Cl_{\Gamma}(M)}L^{2}(\Lambda^{\ast}(M)),Q=N_{1}(\sigma_{tcl}^{\prime}(P)\hat{\otimes}1)+N_{2}(1\otimes\widetilde{\mathcal{D}_{M,\tilde{\Gamma}}^{cl}})),$
    where $E_{1}$ is the Hilbert module of the sections of the family of Hilbert spaces $\{E_{x}\hat{\otimes}Cl_{\tilde{\Gamma}_{x}}(M)\otimes L^{2}(T_{x}M)\}_{x\in M}$, $N_{1},N_{2}$ are pseudo-differential operators constructed in the $KK$-product, and $\widetilde{\mathcal{D}_{M,\tilde{\Gamma}}^{cl}}$ is the $KK$-theoretic connection for $\mathcal{D}_{M,\tilde{\Gamma}}^{cl}$. Note that the Clifford terms act over the space $\Lambda^{\ast}(TM)$.
  • (iii)
    Let $\xi$ be the multiplication variable on $TM$, $\chi$ and $\zeta$ be the derivation variables. The operator $\sigma^{\prime}(P)$ and the symbol term $f_{\Gamma}$ take variables in $x$ and $\xi$, and the symbol of $\mathcal{D}_{M,\tilde{\Gamma}}^{cl}$ takes variables in $\chi$ and $\zeta$. Our rotation homotopy is defined using the symbol of $\sigma^{\prime}(P)\otimes 1$ and $\widetilde{\mathcal{D}^{cl}_{M,\tilde{\Gamma}}}$ by rotation of variables
    $\xi\mapsto\cos t\ \xi-\sin t\ \chi;\ \chi\mapsto\sin t\ \xi+\cos t\ \chi,\ t\in[0,\pi/2].$
    Note that $f_{\Gamma}$ depends on the orbits whose dimensions may jump at some points, so we construct the rotation homotopy to be induced from the rotation homotopy of the $\mathfrak{b}$ symbol of the orbital Dirac operator $\mathfrak{F}_{G}$, defined in Definition 5.5 in [kasparov2022k], on the trivial vector bundle $\mathfrak{g}_{M}\cong M\times\mathfrak{g}$. The variables corresponding to $\xi,\chi$ for the $\mathfrak{b}$ symbol are $\eta$ and $\theta$. The rotation homotopy for $f_{\Gamma}$ is taken between $\eta$ and $\theta$.
  • (iv)
    Such a homotopy involves both the multiplication variable and derivation variable, so it is necessary to check that the commutator of each two terms is 00 in their corresponding $KK$-classes, so that the homotopy can be well-defined in the operator sense. The only non-trivial case we need to consider is that $f_{\Gamma}$ does not commute well with the operator $\widetilde{\mathcal{D}^{cl}_{M,\tilde{\Gamma}}}$ because these two operators use the same Clifford variables. However, the algebra $C^{\ast}(G,C(M))$ contains the multiplier $(1+\Delta_{G})^{-1}$ which kills the non-trivial commutators. Along this rotation homotopy, the commutator $[f_{\Gamma},\widetilde{\mathcal{D}^{cl}_{M,\tilde{\Gamma}}}]$ will not be killed by the multiplication with $C^{\ast}(G,C(M))$. However, by the important remark in the proof of Theorem 10.3 of [kasparov2022k], the homotopy can be corrected to have norm $1$ and at the end of the homotopy, the non-trivial commutator again vanishes. As a result, we get a well-defined rotation homotopy as desired.
  • (v)
    Use functional calculus on the operator $P$ such that $\lVert P\rVert\leq 1$. Consider the operator $T=P\hat{\otimes}1+(1-{P}^{2})^{1/2}\hat{\otimes}\mathfrak{F}_{G}$ corresponding to the symbol class $[\sigma^{tcl}_{H}(P)]\otimes_{C^{\ast}(Gr(H))}[\mathcal{E}]$ act on $L^{2}(E_{1}\hat{\otimes}\Lambda^{\ast}(TM))$. It suffices to show that $1-T^{2}$ is a compact operator, so that such an operator $T$ provides a representation of a $KK$-class. As in the proof of Theorem 10.3 of [kasparov2022k], the commutator term $[P\hat{\otimes}1,1\hat{\otimes}\mathfrak{F}_{G}]$ is estimated on two sides by the operator $(1+\mathcal{D}_{O}^{2})^{-1/2}$, i.e., for some positive constant number $c$, we have
    $-c(1+\mathcal{D}_{O}^{2})^{-1/2}\leq[P\hat{\otimes}1,1\hat{\otimes}\mathfrak{F}_{G}]\leq c(1+\mathcal{D}_{O}^{2})^{-1/2}.$
    By calculation, $1-T^{2}=(1-P^{2})(1-\mathfrak{F}_{G}^{2})-(1-P^{2})^{1/2}[P\hat{\otimes}1,1\hat{\otimes}\mathfrak{F}_{G}]$. Note that by definition, the transversally H-elliptic operator $P$ satisfies that $(1-\sigma_{H}(P)^{2})(1+\sigma_{H}(\Delta_{G}))^{-1}$ is compact and thus has negative H-order. So by Theorem 1.6.13, $(1-P^{2})(1+\mathcal{D}_{O}^{2})^{-1}$ is a compact operator. Thus $1-T^{2}$ is also compact.
    The operator $\sigma^{\prime}(P)(x,\xi)$ is defined on $TM$ and the operator $P$ can be lifted to $TM$ with the same symbol, to be a $KK$-theoretic connection $\tilde{P}$. Note that $\sigma^{\prime}(P_{0})(x,\xi)$ is the Fourier transform of the freezing coefficient of the H-principal part of $P_{0}$, so the rotation homotopy transforms the multiplication with $\xi$-variable by the differential by $x$-variable in the operator sense. $\sigma^{\prime}(P_{0})(x,\xi)$ on $TM$ is a polynomial in $\xi$ variable, so it is transformed to an operator having the same $H$-principal part with $\tilde{P_{0}}$ on $TM$. Thus by functional calculus, $\sigma^{\prime}(P)(x,\xi)$ on $TM$ is transformed to $\tilde{P}$ up to lower order terms. As a result, we see that the lift of $T=P\hat{\otimes}1+(1-{P}^{2})^{1/2}\hat{\otimes}\mathfrak{F}_{G}$ to $TM$ gives a representation of the operator corresponding to the symbol $\sigma^{\prime}(P)(x,\chi)\hat{\otimes}1+(1-\sigma^{\prime}(P)^{2}(x,\chi))^{1/2}\hat{\otimes}f_{{\Gamma}}(x,\chi)$ after the rotation homotopy on $\sigma^{\prime}_{tcl}(P)$.
    As in the proof of Theorem 3.2.6, by executing the rotation homotopy, the Dolbeault component $\mathcal{D}_{M,\tilde{\Gamma}}^{cl}$ in $Q$ is transitioned into the assembly of Bott–Dirac operators $\mathcal{C}=\{\mathcal{C}_{x},x\in M\}$ on $L^{2}(\Lambda^{\ast}(T_{x}M))$.
    By the statements above, $Q$ can be transformed by the rotation homotopy into the operator $N_{1}\tilde{T}+N_{2}1\hat{\otimes}\mathcal{C}$, representing the $KK$-product of $[T]$ and $[\mathcal{C}]$. The operator $\mathfrak{F}_{G}$ in $T$ can be removed since it is killed by the multiplication with $C(M)\cdot(1+\Delta_{G})^{-1}\subset C^{\ast}(G,C(M))$. Note also that the Bott-Dirac operators $\mathcal{C}$ correspond to the $1$ element in $KK(C^{\ast}(G,C(M)),C^{\ast}(G,C(M)))$. As a result, after the rotation, we arrive at the index class $[P]\in KK(C^{\ast}(G,C(M)))$.
  • (vi)
    Consider the induced restriction from $KK^{0}(C^{\ast}(G,C(M)))$ to $KK^{0}(C^{\ast}(G))$ and from $KK^{G}(C(M),C^{\ast}(Gr(H))\otimes_{C(M)}Cl_{\tilde{\Gamma}}(M))$ to $KK^{G}(\mathbb{C},C^{\ast}(Gr(H))\otimes_{C(M)}Cl_{\tilde{\Gamma}}(M))$ by the inclusion $\mathbb{C}\subset C(M)$, we will get analytical index class $[P]$ which is the left-hand side.
That finishes the proof.
Remark 5.3.1.
It is considerable to establish the index theory for transversally H-elliptic operators via an adiabatic groupoid approach, similar to the construction in Section 3.1. For each representation of $G$, it entails contemplating distinct $C^{\ast}$-algebras for each fiber of the adiabatic groupoid, so that it can represent a commutative diagram and we obtain an index in $\mathbb{Z}$. By such construction, the $C^{\ast}$-algebra diagram induces an element in $\text{Hom}(R(G),\mathbb{Z})\cong\widehat{R(G)}\cong K^{0}(C^{\ast}(G))$, and it should be consistent with the analytical index class of $P$. A different proof of the index theorem using such a groupoid approach is left for future research.
Chapter 6 Verification of the condition and basic examples
6.1 Fourier transform of nilpotent Lie group $C^{\ast}$-algebras
In this chapter, we focus on verifying the compactness condition outlined in Definition 4.5.1, i.e., let $e$ be an element in the multiplier algebra of the $C^{\ast}(Gr(H)_{x})$-module $E=C^{\ast}(Gr(H)_{x})$, $e\in M(C^{\ast}(Gr(H)_{x}))$, we want to show whether or not $e$ lies in $\mathcal{K}(E)=E$.
Fourier transform from ordinary cosymbol to symbol
Recall that for a differential operator $P$, let $F$ be the ordinary cosymbol of $P$ we discussed in Remark 1.4.6. We can take the Fourier transformation of $F\in M(C^{\ast}(TM))$ to the symbol $f=\widehat{F}$ in $M(C_{0}(T^{\ast}M))=C_{b}(T^{\ast}M)$. Let us describe an idea to show the compactness of $F$ roughly as follows: We show that, when $\xi$ converges to infinity, $f(x,\xi)$ converges to 00, and thus, it lies in $C_{0}(T^{\ast}M)$ and then the original cosymbol operator $F$ is a compact multiplier of $C^{\ast}(TM)$. See the commutative diagram below:
[…]
In case of a differential operator $P$ on the vector bundle $E$, the ordinary symbol can similarly be understood as an element in $C_{b}(M;p^{\ast}E)\cong C_{b}(T^{\ast}M)\otimes_{C(M)}C(M;E)$, with a compact symbol in the subspace $C(M;p^{\ast}E)\cong C_{0}(T^{\ast}M)\otimes_{C(M)}C(M;E)$.
Main idea to verify the compactness condition
For an element $e\in M(C^{\ast}(\mathcal{G}))$, we show its Fourier transform $\mathcal{F}(e)\in\mathcal{F}(M(C^{\ast}(\mathcal{G})))$ lies in $\mathcal{F}(C^{\ast}(\mathcal{G}))$ by using the $C^{\ast}$-structure of the resulting space of Fourier transform. Then, $e\in\mathcal{K}(C^{\ast}(\mathcal{G}))=C^{\ast}(\mathcal{G})$. A parallel diagram illustrates the relationship:
[…]
We can apply $\mathcal{G}$ to be $Gr(H)_{x}$ for each $x$, which is nilpotent. In this case, we can verify the compactness of a multiplier on $C^{\ast}(\mathcal{G})$ for a nilpotent Lie group $\mathcal{G}$ by using the theory regarding Fourier transform of nilpotent Lie group $C^{\ast}$-algebra, which is treated in [beltictua2017fourier].
Fourier transform of nilpotent Lie group $C^{\ast}$-algebras
First, We introduce some definitions given in [beltictua2017fourier].
Definition 6.1.1.
Let $A$ be a $C^{\ast}$-algebra with spectrum $\widehat{A}$, which consists of the unitary equivalence classes of irreducible representations. We choose for every $\gamma\in\widehat{A}$ a representation $\pi_{\gamma}:A\to B(\mathcal{H}_{\gamma})$ in the equivalence class $\gamma$.
Define $\mathit{l}^{\infty}(\widehat{A})$ to be the algebra of all bounded operator fields defined over $\widehat{A}$ by
$\mathit{l}^{\infty}(\widehat{A}):=\big{\{}\phi=(\phi({\gamma}))\in\mathcal{B}(\mathcal{H}_{\gamma})_{\gamma\in\widehat{A}}\ |\ \lVert\phi\rVert=\sup_{\gamma}\ \lVert\phi(\gamma)\rVert_{B(\mathcal{H}_{\gamma})}<\infty\big{\}}.$
Then we define for $a\in A$ its Fourier transform $\mathcal{F}_{A}(a)=\widehat{a}\in\mathit{l}^{\infty}(\widehat{A})$ by
$\mathcal{F}_{A}(a)(\gamma)=\widehat{a}(\gamma)=\pi_{\gamma}(a).$
Definition 6.1.2.
  • (i)
    Let $S$ be a topological space. We say that $S$ is locally compact of step $\leq d$ if there exists a finite increasing family
    $\emptyset\neq S_{d}\subset S_{d-1}\subset\cdots\subset S_{0}=S$
    of closed subsets of $S$, such that the subsets $\Gamma_{d}=S_{d}$, $\Gamma_{i}=S_{i-1}\setminus S_{i}$ for $11\leq i\leq d-$ 1 are locally compact and Hausdorff in their relative topologies.
  • (ii)
    Let $S$ be locally compact of step $\leq d$, and let $\{\mathcal{H}_{i}\}_{{i=1,\dots,d}}$ be Hilbert spaces. For a closed subset $M\subset S$, denote by $CB(M)$ the unital $C^{\ast}$-algebra of all uniformly bounded operator fields $(\psi(\gamma)\in\mathcal{B}(\mathcal{H}_{i}))_{\gamma\in S\cap\Gamma_{i},i=1,\dots,d}$, which are operator norm continuous on the subsets $\Gamma_{i}\cap M$ for every $i\in\{1,\dots,d\}$ with $\Gamma_{i}\cap M\neq\emptyset$. We provide the algebra $CB(M)$ with the infinity-norm
    $\lVert\phi\rVert_{M}:=\sup\big{\{}\lVert\phi(\gamma)\rVert_{\mathcal{B}(\mathcal{H}_{i})}\ |\ M\cap\Gamma_{i}\neq 0,\ \gamma\in M\cap\Gamma_{i}\big{\}}.$
Definition 6.1.3.
We say that the $C^{\ast}$-algebra $A$ has norm-controlled dual limits if the spectrum $\widehat{A}$ of $A$ is a locally compact space of step $\leq d$, and for every $a\in A$, we have
  • (1)
    The mappings $\gamma\to\mathcal{F}(a)(\gamma)$ are norm continuous on the difference sets $\Gamma_{i}=S_{i-1}\setminus S_{i}$;
  • (2)
    For any $i=0,\dots,d+1$ and any converging sequence contained in $\Gamma_{i}$ with limit set outside $\Gamma_{i}$, thus in $S_{i}$, there exists a properly converging subsequence $\overline{\gamma}=(\gamma_{k})$, $k\in\mathbb{N}$, a constant $C>0$ and for every $k\in N$ an involutive linear mapping $\sigma_{\overline{\gamma},k}:CB(S_{i})\to B(\mathcal{H}_{i})$, bounded by $C\lVert\cdot\rVert_{S_{i}}$, such that for arbitrary $a\in A$,
    $\lim_{k\to\infty}\lVert F(a)(\gamma_{k})-\sigma_{\overline{\gamma},k}(\mathcal{F}(a)|_{S_{i}})\rVert_{B(\mathcal{H}_{i})}=0.$
Theorem 6.1.4 ([beltictua2017fourier]).
For a connected, simply connected nilpotent Lie group $\mathcal{G}$, the $C^{\ast}$-algebra $C^{\ast}(\mathcal{G})$ has norm-controlled dual limits.
Definition 6.1.5.
Let $S$ be a locally compact topological space of step $\leq d$. Define $B^{\ast}(S)$ to be the set of all operator fields $\phi$ defined over $S$ such that
  • (a)
    $\phi(\gamma)\in K(\mathcal{H}_{i})$ for every $\gamma\in\Gamma_{i}=S_{i-1}\setminus S_{i},i=1,\dots,d$;
  • (b)
    The field $\phi$ is uniformly bounded, i.e.,
    $\lVert\phi\rVert=\max_{i}\sup_{\gamma\in\Gamma_{i}}\lVert\phi(\gamma)\rVert_{B(\mathcal{H}_{i})}<\infty;$
  • (c)
    The mappings $\gamma\mapsto\phi(\gamma)$ are norm continuous on the difference sets $\Gamma_{i}$;
  • (d)
    For any sequence $(\gamma_{k})_{k\in\mathbb{N}}\subset S$ going to infinity,
    $\lim_{k\to\infty}\lVert\phi(\gamma_{h})\rVert_{\text{op}}=0;$
  • (e)
    For any $i=0,\dots,d+1$ and any converging sequence contained in $\Gamma_{i}$ with its limit set outside $\Gamma_{i}$, in $S_{i}$, there exists a properly converging subsequence $\overline{\gamma}=(\gamma_{k})$, $k\in\mathbb{N}$, a constant $C>0$ and for every $k\in N$ an involutive linear mapping $\sigma_{\overline{\gamma},k}:CB(S_{i})\to B(\mathcal{H}_{i})$, which is bounded by $C\lVert\cdot\rVert_{S_{i}}$, such that for arbitrary $a\in A$,
    $\lim_{k\to\infty}\lVert\phi(\gamma_{k})-\sigma_{\overline{\gamma},k}(\phi|_{S_{i}})\rVert_{B(\mathcal{H}_{i})}=0.$
Theorem 6.1.6 ([beltictua2017fourier]).
Let $A$ be a $C^{\ast}$-algebra with norm-controlled dual limits. Then, the Fourier transform of $A$ can be derived as the $C^{\ast}$-algebra $B^{\ast}(\widehat{A})$.
Consequently, combining Theorem 6.1.4 and Theorem 6.1.6, we have
Corollary 6.1.7.
The Fourier transform of $C^{\ast}(\mathcal{G})$ is the $C^{\ast}$-algebra $B^{\ast}(\widehat{\mathcal{G}})$.
Example 6.1.8.
The Heisenberg group $C^{\ast}$-algebra $C^{\ast}(\mathbb{H}_{n})$ is isomorphic to $D_{\nu}(\widehat{\mathbb{H}_{n}})$ by Fourier transform, where $D_{\nu}(\widehat{\mathbb{H}_{n}})$ is defined to be the algebra comprising all operator fields $(F=F(\lambda))_{\lambda\in\mathbb{R}}$ satisfying the following conditions:
  • $F(\lambda)$ denotes a compact operator on $L^{2}(\mathbb{R}^{2})$;
  • $F(0)\in C^{\ast}(\mathbb{R}^{2})$;
  • the mapping $\mathbb{R}^{\ast}\to B(L^{2}(\mathbb{R}^{n}))):\lambda\to F(\lambda)$ is norm continuous;
  • $\lim_{\lambda\to\infty}\lVert F(\lambda)\rVert_{op}=0$;
  • $\lim_{\lambda\to 0}\lVert F(\lambda)-\nu_{\lambda}(F(0))\rVert_{op}=0$, where
[…]
    where $\eta$ denotes a Schwartz-function in $S(\mathbb{R}^{n})$ with $L^{2}$-norm equal to $1$, $P_{\eta}$ is the projection onto the one-dimensional subspace $\mathbb{C}\eta$.
In this case, $C^{\ast}(\mathbb{H}_{n})$ of $\mathbb{H}_{n}$ is an extension of an ideal $J$ isomorphic to $C_{0}(\mathbb{R}^{\ast},\mathcal{K})$ with the quotient algebra isomorphic to $C^{\ast}(\mathbb{R}^{2n})$.
6.2 Application to the multiplier algebra
The multiplier algebra of $C_{0}$ functions on a topological space $X$ with values in a certain $C^{\ast}$-algebra can be calculated as follows.
Definition 6.2.1.
Let $A$ be a $C^{\ast}$-algebra. Let $M(A)$ be its multiplier $C^{\ast}$-algebra. The strict topology on $M(A)$ is given by the convergence condition:
$x_{\lambda}\to x\Longleftrightarrow\forall a\in A:(\lVert xa_{\lambda}-xa\rVert+\lVert a_{\lambda}x-ax\rVert\to 0).$
Lemma 6.2.2 ([akemann1973multipliers] Corollary 3.4).
Let $A$ be a $C^{\ast}$-algebra, then
${M}(C_{0}(X,A))=C_{b}(X,M(A)_{\beta}),$
where $M(A)_{\beta}$ is the multiplier algebra of $A$ with the strict topology.
Remark 6.2.3.
By Lemma 6.2.2, we see that the multiplier algebra for continuous functions valued in compact operator space consists of continuous elements valued in the bounded operator space with the strict topology.
Applying Remark 6.2.3 to the conditions in Definition 6.1.5, we get
Corollary 6.2.4.
A multiplier of $M(B^{\ast}(S))$ is an element of the operator field $\psi$ over $S$ that satisfies:
  • $(a^{\prime})$
    $\psi(\gamma)\in B(\mathcal{H}_{i})$ for every $\gamma\in\Gamma_{i}=S_{i-1}\setminus S_{i},i=1,\dots,d$;
  • $(b^{\prime})$
    The field $\psi$ is uniformly bounded;
  • $(c^{\prime})$
    The mappings $\gamma\mapsto\psi(\gamma)$ are strictly continuous on the difference sets $\Gamma_{i}$;
  • $(d^{\prime})$
    For any $i=0,\dots,d+1$ and any converging sequence contained in $\Gamma_{i}$ with limit set outside $\Gamma_{i}$, thus in $S_{i}$, there exists a properly converging subsequence $\overline{\gamma}=(\gamma_{k})$, $k\in\mathbb{N}$, a constant $C>0$ and for every $k\in N$ an involutive linear mapping $\sigma_{\overline{\gamma},k}:CB(S_{i})\to B(\mathcal{H}_{i})$, which is bounded by $C\lVert\cdot\rVert_{S_{i}}$, such that for arbitrary $a\in A$,
    $\lim_{k\to\infty}\psi(\gamma_{k})-\sigma_{\overline{\gamma},k}(\psi|_{S_{i}})=0$
    in the strict topology of $B(\mathcal{H}_{i})$.
Take a multiplier $F\in M(C^{\ast}(\mathcal{G}))$, $F:C^{\ast}(\mathcal{G})\to C^{\ast}(\mathcal{G})$. For each representation $\pi\in\widehat{\mathcal{G}}$, $\pi$ extends to a representation of $C^{\ast}(\mathcal{G})$; it induces a map $\pi(F):=F\otimes_{\pi}1$ from $C^{\ast}(\mathcal{G})\otimes_{\pi}\mathcal{H}_{\pi}\cong\mathcal{H_{\pi}}$ to itself. Then by Definition 6.1.1, we find that $\{\pi(F)\}_{\pi\in\widehat{\mathcal{G}}}$ is the Fourier transform of an element $F$ in the multiplier algebra $M(C^{\ast}(\mathcal{G}))$.
Corollary 6.2.5.
A multiplier $Q\in M(E_{x}\otimes C^{\ast}(Gr(H)_{x}))$ is compact, i.e., $Q\in\text{End}(E_{x})\otimes C^{\ast}(Gr(H))_{x}$ if and only if the operator field
$\phi=\{\pi_{\gamma}(Q)\}_{\gamma\in\widehat{Gr(H)}_{x}}\subseteq\text{End}(E_{x})\otimes\mathit{l}^{\infty}(\widehat{Gr(H)}_{x})$
lies in the subspace $\text{End}(E_{x})\otimes B^{\ast}(\widehat{Gr(H)}_{x})$, i.e., each term $\phi_{ij}$ of the $\mathit{l}^{\infty}(\widehat{Gr(H)}_{x})$-valued matrix $\phi$ in $\text{End}(E_{x})$ satisfies the conditions $(a)-(e)$ in Definition 6.1.5.
Proof.
Take $S=C^{\ast}(\mathcal{G})$ and where $G=Gr(H)_{x}$. The multiplier $Q$ lies in $\text{End}(E_{x})\otimes C^{\ast}(Gr(H))_{x}$ if and only if its Fourier transform $\phi=\{\pi_{\gamma}(Q)\}_{\gamma\in\widehat{Gr(H)}_{x}}$ lies in $\text{End}(E_{x})\otimes\mathcal{F}(C^{\ast}(Gr(H)_{x}))$ which is equal to $\text{End}(E_{x})\otimes B^{\ast}(\widehat{Gr(H)}_{x})$ by Corollary 6.1.7. By Definition 6.1.5, it is equivalent to the conditions $(a)-(e)$ are satisfied for each term $\phi_{ij}$ of the $\mathit{l}^{\infty}(\widehat{Gr(H)}_{x})$-valued matrix $\phi$ in $\text{End}(E_{x})$. ∎
Example 6.2.6.
For instance, consider the $\mathbb{R}^{2}$ action on $\mathbb{H}_{3}\cong\mathbb{R}^{5}$ on the first two $x_{1},y_{1}$ directions. In addition, consider the operators of the form $P_{\gamma}=-X_{2}^{2}-Y_{2}^{2}+i\gamma T$, which is induced from an operator on $\mathbb{H}_{2}\cong\mathbb{R}^{3}$ to $\mathbb{H}_{3}\cong\mathbb{R}^{5}$. $P$ is normalized to be self-adjoint, and $\phi$ is calculated as the diagonal matrix with
$\phi_{11}=(1+\sigma_{H}(P^{\ast},x)\sigma_{H}(P,x))^{-1}(1+d/dx_{1}^{2}+d/dy_{1}^{2})^{-1},$
$\phi_{22}=(1+\sigma_{H}(P,x)\sigma_{H}(P^{\ast},x))^{-1}(1+d/dx_{1}^{2}+d/dy_{1}^{2})^{-1}.$
Suppose $P_{\gamma}$ is an H-elliptic operator on $\mathbb{R}^{3}$, i.e., $\gamma$ is not an odd integer, then $(b)(c)(e)$ will be evident by the condition that each representation of $\mathbb{H}_{3}$ can be pulled back to the representation of $\mathbb{R}^{2}$ and $\mathbb{H}_{2}\cong\mathbb{R}^{3}$—both of which satisfies these conditions. We verify the conditions $(a)$ and $(d)$. Based on Example 2.2, the representations of $\mathbb{H}_{3}$ contain two classes. For $(a)$, using the coadjoint orbit, we can derive:
  • for a Schrödinger representation $\pi_{\lambda}$ of $\mathbb{H}_{3}$ which can be represented by $\lambda T$, the restriction to $\mathbb{H}_{2}$ will be Schrödinger representations $\pi^{\prime}_{\lambda}$ of $\mathbb{H}_{2}$. Based on the Fourier transform of the sub-Lie group $\mathbb{H}_{2}\subset\mathbb{H}_{3}$, $\pi^{\prime}_{\lambda}(1+\sigma_{H}(P^{\ast},x)\sigma_{H}(P,x))^{-1})$ and $\pi^{\prime}_{\lambda}(1+\sigma_{H}(P,x)\sigma_{H}(P^{\ast},x))^{-1})$ will be compact on $L^{2}(\mathbb{R}^{1})$ regarding as a representation of an element of the liminary $C^{\ast}$-algebra $C^{\ast}(\mathbb{H}_{2})$, generated by $x_{2},y_{2},t$ coefficient. $\pi_{\lambda}((1+d/dx_{1}^{2}+d/dy_{1}^{2})^{-1})$ will be compact on the first variable $L^{2}(\mathbb{R}^{1})$, as it can be regarded as a representation of an element of the liminary $C^{\ast}$-algebra $C^{\ast}(\mathbb{H}_{2})$, generated by $x_{1},y_{1},t$ coefficient. Upon combining these two parts, we observe that $\pi_{\lambda}(\phi)$ is compact for all Schrödinger representations $\pi_{\lambda}$.
  • for one-dimensional representations, if $\pi=\pi_{x_{1},x_{2},y_{1},y_{2}}$ is presumed to be nonzero for the $x_{1},x_{2}$ components, then $\pi((1+d/dx_{1}^{2}+d/dy_{1}^{2})^{-1})$ will be compact on $\mathbb{R}$. Otherwise, if it is nonzero for the $y_{1},y_{2}$ part, then $\pi(1+\sigma_{H}(P^{\ast},x)\sigma_{H}(P,x))^{-1})$ and $\pi(1+\sigma_{H}(P,x)\sigma_{H}(P^{\ast},x))^{-1})$ will be compact on $\mathbb{R}$. Consequently, the product will be a compact operator in all cases.
$(d)$ can also be shown by the same argument. As a result, $P_{\gamma}$ is a transversally H-elliptic operator on $\mathbb{H}_{3}\cong\mathbb{R}^{5}$ if $\gamma$ doesn’t take values in odd integers.
Remark 6.2.7.
  • (i)
    In general cases, the direct decomposition of representations (aligned with coadjoint orbits) into transverse and leaf-wise components, similar to the partitioning of cotangent vectors, presents significant challenges. This complexity leads to the difficulties involved in addressing transversal H-ellipticity. Nevertheless, the application of Fourier transform methodologies facilitates the transformation of operators to actions on Hilbert spaces, thereby simplifying the task to prove the compactness of these operators within each Hilbert space. This approach allows for the partitioning of representing Hilbert spaces into distinct segments, facilitating a separate examination of the transverse and leaf-wise directional elements.
  • (ii)
    If we focus on the symbol terms of differential operators and the inverses of Rockland terms within the cosymbol space, it is possible that the continuity conditions $(c)$ and $(e)$ in Definition 6.1.5 are redundant. However, the absence of a definitive proof leaves this aspect open to further research.
Appendix A Appendix
A.1 Other ideas for the definition of transversal H-ellipticity
Alternative definitions of transversal H-ellipticity can be considered. All the definitions are motivated by some characteristics of H-ellipticity or transversal ellipticity. However, these definitions are not sufficient to calculate in $KK$-theory and it’s hard to show the equivalent or inclusion relations between any two of these different conditions, as the $C^{\ast}$-algebra $C^{\ast}(Gr(H))$ is non-commutative and we cannot regularly decompose $gr(H)_{x}$ elements into transversal and leaf-wise directions. These definitions only provide other possible directions to analyze.
We will briefly show the possible directions for analyzing operators in these different definitions.
A.1.1 The main definition
Recall that our main definition is
Definition A.
Let $M$ be a compact filtered Riemannian manifold and $G$ be a Lie group acting on $M$ isometrically and preserving the filtered structure. We call a pseudodifferential operator of 00-order $P$ is transversally H-elliptic if at each point $x\in M$, the multiplier term
$(1-\sigma_{H}^{2}(P,x))(1+\sigma_{H}(\Delta_{G},x))^{-1}\in M(E\otimes_{C(M)}C^{\ast}(Gr(H)_{x}))$
is compact as operators on $C^{\ast}(Gr(H)_{x})$-modules.
The main definition (Definition A) is motivated by the Rockland condition characteristic of H-elliptic operators and the tangent Clifford symbol for transversally elliptic operators. It facilitates the definition of index and symbol classes in $KK$-theory for operator $P$, enabling the application of Kasparov’s method for index theorem analysis.
A.1.2 Definition from Atiyah’s approach
The second definition is stated as follows:
Definition B.
$P$ is called transversally H-elliptic if for each $x\in M$, for all representation $\pi$ in $Gr(H)_{x}$, both $\pi(\sigma_{H}((P,\Delta_{G}),x))$ and $\pi(\sigma_{H}((P^{\ast},\Delta_{G}),x))$ are injective.
Inspired by Atiyah’s approach [atiyah2006elliptic], the second definition (Definition B) focuses on the injectivity of symbols for the operator $P$ and its adjoint $P^{\ast}$ when combined with a leaf-wise elliptic Laplacian operator. This perspective allows for the analytical index to be defined as a distribution over the group $G$, employing Atiyah’s method.
More explicitly, let
$H^{s}(E)_{\lambda}=\ker(\Delta_{G}-\lambda):H^{s}(E)\to H^{s-2}(E),$
where $H^{s}(E)$ are ordinary Sobolev spaces and $\Delta_{G}$ is the orbital Laplacian operator we defined in Definition 4.1.3. Since $P$ commutes with the $G$-action, it commutes with $\Delta_{G}$, and especially, it commutes the $\lambda$-eigenspaces $\mathcal{D}(E)_{\lambda}\subset\mathcal{D}(E)$ and $\mathcal{D}(F)_{\lambda}\subset\mathcal{D}(F)$, where $\mathcal{D}(E)=C^{\infty}_{c}(E)$. We obtain a linear map $P_{\lambda}:\mathcal{D}(E)_{\lambda}\to\mathcal{D}(F)_{\lambda}$ and it extends to
$P_{\lambda}:H^{s}(E)_{\lambda}\to H^{s-k}(F)_{\lambda}.$
Similarly, we can construct the operator $P_{\lambda}^{\ast}$ from $P^{\ast}$. Based on the condition in (Definition B), such $P_{\lambda}$ and $P^{\ast}_{\lambda}$ are Fredholm for each $\lambda$. Its index is defined as
${Ind}(P_{\lambda})=\chi_{\ker(P_{\lambda})}-\chi_{\ker(P^{\ast}_{\lambda})},$
where $\chi_{\ker(P_{\lambda})}$ and $\chi_{\ker(P^{\ast}_{\lambda})}$ are defined to be distributions over $G$. The analytical index of $P$ is then defined as
${Ind}(P)=\sum_{\lambda}{Ind}(P_{\lambda})$
as an equality of distributions on $G$. It will be possible that for operators in Definition B, we can define the cohomological index and prove the index theorem that the cohomological index is equal to the analytical index.
However, the $\Delta_{G}$ operator cannot be killed by certain special $\pi_{\xi}$. So for an operator $P$ satisfying the condition in Definition B, it is difficult to separate the $(P,\Delta_{G})$ or $(P^{\ast},\Delta_{G})$ in the condition to obtain the operator $P$ itself. Therefore, it is hard to define the symbol class and index class without $\Delta_{G}$ in $KK$-homology. The analysis of operators in this definition in $KK$-theory seems extremely difficult.
A.1.3 An intuitive definition
In Chapter 4, we have defined the bundle isomorphism $\psi:TM\to gr(H)$, which maps a vector $v\in T_{x}M$ decomposed as $v=\sum_{i}v^{i}$ to $\psi(v)=\sum_{i}[v^{i}]\in gr(H)_{x}$, with each $v^{i}\in H^{i}_{x}$ and all $v^{i}$’s are orthogonal to each other. We obtain the adjoint map $\psi^{\ast}:gr(H)^{\ast}\to T^{\ast}M$. The image of transversal part $T^{\ast}_{G,x}M$ by $(\psi^{\ast})^{-1}$ is denoted as $U^{\ast}_{G,x}\subseteq gr(H)_{x}^{\ast}$. We define
$O^{\ast}_{G,x}=\{\xi\in gr(H)^{\ast}_{x}\ |\text{ The coadjoint orbit containing }\xi\text{ passing through }U^{\ast}_{G,x}\}.$
It can be understood as the combination of all the orbits containing a point in $U_{G,x}^{\ast}$.
The third definition is stated as follows:
Definition C.
$P$ is called transversally H-elliptic if for all $x\in M$ $\sigma(P,x,\xi)$ is bijective for $\pi\in O^{\ast}_{G,x}$, i.e., $\sigma(P,x,\pi_{\xi})$ is bijective if the coadjoint orbit corresponding to $\pi_{\xi}$ contains at least one point of $T^{\ast}_{G}M$.
The third definition (Definition C) is the most straightforward one, directly aligning with the foundational concept of transversal ellipticity. However, its practical application within $KK$-theory is restricted by the lack of information on the representations out of $O^{\ast}_{G,x}$. There is also a problem that we cannot decompose coadjoint orbits of the nilpotent Lie group $Gr(H)_{x}$, similar to the decomposition of $T_{x}M$ to tangent part and transversal part, because $C^{\ast}(Gr(H))$ is non-commutative,
Adopting this third definition (Definition C) facilitates a construction parallel to that for transversally elliptic cases by defining a specialized $KK$-class
$[\mathcal{E}^{\prime}]\in KK_{G}(C^{\ast}(O_{G}),C_{0}(T^{\ast}_{G}M)).$
Define
$C^{\ast}(O_{G,x})=\frac{C^{\ast}(Gr(H)_{x})}{\bigcap_{\xi\in O_{G,x}}\ker\pi_{\xi}}.$
The $KK$-class $[\mathcal{E}]$ can be restricted to define a $KK$-class $[\mathcal{E}^{\prime}]\in KK_{G}(C^{\ast}(O_{G}),C_{0}(T^{\ast}_{G}M))$.
Thus, if we consider such transversally H-elliptic operators in Definition C with “good” symbol (which is defined in [berline1996chern]), we can typically define the symbol class $[\sigma^{\prime}_{H}(P)]\in KK_{G}(\mathbb{C},$ $C^{\ast}(O_{G}))$, and index class $[P]\in KK(C^{\ast}(G),\mathbb{C})$. Then $[\mathcal{E}^{\prime}]$ in a certain sense yields a restricted Connes-Thom isomorphism. If we define the Dolbeault operator $[\mathcal{D}_{M}^{\prime}]\in K^{0}(C^{\ast}(G,C_{0}(T^{\ast}_{G}M)))$ in transversally direction similarly, we obtain a formula that
$[P]=j^{G}([\sigma^{\prime}_{H}(P)]\otimes_{C^{\ast}(O_{G})}[\mathcal{E}^{\prime}])\otimes_{C^{\ast}(G,C_{0}(T^{\ast}_{G}M))}[\mathcal{D}_{M}^{\prime}].$
[…]